### v1

# A Neural-Guided Variational Quantum Algorithm for Efficient Sign Structure Learning in Hybrid Architectures

Abstract
Variational quantum algorithms hold great promise for unlocking the power of near-term quantum processors, yet they are frequently hindered by high measurement costs, barren plateaus, and challenging optimization landscapes. Here, we introduce sVQNHE, a neural-guided variational quantum algorithm that decouples amplitude and sign learning across classical and quantum modules, respectively. Our approach employs shallow quantum circuits composed of commuting diagonal gates to efficiently model quantum phase information, while a classical neural network learns the amplitude distribution and guides circuit optimization in a bidirectional feedback loop. This hybrid quantum-classical synergy not only reduces measurement costs but also achieves high expressivity with limited quantum resources and improves the convergence rate of the variational optimization. We demonstrate the advancements brought by sVQNHE through extensive numerical experiments (including realistic noisy simulations and sampling-based simulations). For the 6-qubit J1-J2 model — a prototypical system with a severe sign problem for Monte Carlo-based methods — it reduces the mean absolute error by $98.9\%$ and suppresses variance by $99.6\%$ relative to a baseline neural network, while requiring nearly $19x$ fewer optimization steps than a standard hardware-efficient VQE. Furthermore, for MaxCut problems on 45-vertex Erdős–Rényi graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$. Importantly, this framework is designed to be scalable and robust against hardware noise and finite-sampling uncertainty, making it well-suited for both current NISQ processors and future high-quality quantum computers. Our results highlight a promising path forward for efficiently tackling complex many-body and combinatorial optimization problems by fully exploiting the synergy between classical and quantum resources in the NISQ era and beyond.
Introduction
Quantum computing promises transformative advances in a variety of applications, from combinatorial optimization to ground-state preparation and dynamical simulation of many-body systems [1, 2]. However, large-scale quantum circuits remain challenging to implement due to hardware limitations, gate errors, and limited qubit coherence [3]. One of the most urgent goals in the noisy intermediate-scale quantum (NISQ) era, and the subsequent early fault-tolerant era [4] (Early FTQC), is to maximize computational utility within these constraints by employing shallow circuits and error-mitigated designs [5].
Among the many strategies proposed to enable practical applications under these conditions, hybrid quantum–classical algorithms, particularly variational quantum algorithms (VQAs) [6, 7, 8], stand out. VQAs combine parameterized quantum circuits (PQC) with classical optimization routines, using quantum circuits as ansatzes to solve problems that can be formulated within a variational framework. Despite initial enthusiasm, standard VQAs face at least three critical challenges. First, optimization challenges, such as barren plateaus [9, 10, 11, 12, 13], where gradients vanish exponentially with circuit depth or system size, severely limit trainability. Second, measurement costs scale poorly with both qubit count and circuit parameters, placing heavy demands on quantum resources. Third, even hardware-efficient circuit designs may require long depths to represent certain wavefunctions, but decoherence makes such computations unreliable in practice [14].
Meanwhile, deep learning has demonstrated remarkable capabilities in learning complex functions and high-dimensional structures from large datasets. In fact, neural networks (NNs) have set new standards for computational efficiency in tasks like quantum simulation [15, 16, 17, 18, 19], which are also commonly associated with quantum computing. A natural question then arises: can quantum resources enhance NN-based variational methods and help push the computational boundary of what is achievable with current and near-future hardware? Previous works have extensively investigated and documented that NNs can efficiently model the wavefunction amplitudes (i.e., essentially the probability distribution), but struggle to capture intricate sign or phase structures in many-body systems [20, 21, 22, 23, 19, 3]. On the other hand, modeling the phase structure of an arbitrary wavefunction can be well approximated with some variations of parameterized Instantaneous Quantum Polynomial (IQP) circuit [24, 25, 26, 27, 28, 29]. This observation motivates one way to combine PQC with NN to jointly model complex wavefunction within a hybrid variational framework. Importantly, such models may exhibit potential quantum advantage if continuous variable IQP (CV-IQP) operations exist for a single iteration [30, 31].
Numerous efforts have explored the combination of quantum computing with deep learning in various settings [32, 33, 34]. An especially relevant example is the variational quantum-neural hybrid eigensolver (VQNHE) [35] algorithm, previously developed by one of us and collaborators. This method uses a neural network to post-process quantum states prepared by a shallow circuit, forming a hybrid quantum-neural ansatz that efficiently expands the accessible quantum state space. However, VQNHE still faces limitations in trainability and incurs non-trivial measurement overhead.
In light of these developments and challenges, we revisit how best to harness both quantum and classical resources within a variational framework, one that maximizes hybrid advantages while mitigating their downsides. To this end, we propose sign-VQNHE (sVQNHE), a hybrid algorithm that comprehensively assuages all the major challenges facing mainstream VQAs: expressiveness, measurement cost, trainability, and robustness (see Table 1). In short, the key innovations of sVQNHE lie in its principled separation of quantum and classical roles: the diagonal layers of the quantum circuit are tasked to model the phase structure (with much reduced measurement costs), while a non-negative neural network only has to learn the amplitude (or effectively, the probability) distribution for improved numerical stability. These components are coupled through an iterative amplitude-transfer mechanism that maps the neural state onto a shallow and classically simulatable quantum layer. Compared to the original VQNHE, this re-design of the hybrid quantum-neural framework not only optimizes expressivity with shallow circuits and substantially improves the measurement costs and the gradient-based convergence rate, but also inherently mitigates sampling noise and decoherence, yielding more robust solutions under realistic conditions (see Results).
During the iterative, bidirectional feedback optimization loop (to train the PQC and NN parameters), if a single diagonal quantum layer cannot fully capture the target wavefunction’s sign structure, we extend the circuit depth by adding a layer of classically simulatable gates along with a new diagonal layer. The classical gate layer approximates the amplitude function learned by the NN by solving a tractable optimization problem, allowing the NN to guide amplitude refinement and assist in layer-wise circuit optimization without encountering barren plateaus. Importantly, because only the parameters of commuting diagonal gates are optimized on the quantum device, the required number of quantum measurements is dramatically reduced compared to the conventional non-commuting circuit ansatzes. In practice, at the last layer of the iterative, bidirectional feedback optimization procedure, one can choose to relax the constraint and allow the neural network to output complex values, thereby enhancing the expressiveness of the hybrid quantum-neural ansatz. Finally, our new algorithm improves the convergence behavior of the original VQNHE by incorporating a stable layer-by-layer training strategy that suppresses numerical instabilities.
This newly proposed iterative and bidirectional feedback loop enables efficient training, expressive state representations, and significant reductions in circuit depth and measurement cost. Our numerical experiments validate these advantages. For instance, in the 6-qubit J1–J2 model, sVQNHE reduces mean absolute error by $98.9\%$ and variance by $99.6\%$ relative to a baseline NN, while achieving nearly 19× faster convergence compared to a hardware-efficient VQE using similar quantum resources. In solving the MaxCut problem on 45-node Erdős–Rényi graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$ over a brickwork ansatz VQE baseline [36].
In summary, sVQNHE offers a scalable and practical path forward for tackling complex many-body quantum simulations and large-scale combinatorial optimization problems by fully leveraging the synergy between classical and quantum computation. As both quantum hardware and neural network models continue to advance, we anticipate that hybrid frameworks like sVQNHE will play an increasingly important role in achieving real-world quantum utility.
Challenge Problem Description sVQNHE Solution
Representation NN-based Variational Monte Carlo struggles to model complex phase structure Delegate the quantum circuit to learn phase structure
Measurement
VQE: $\mathcal{O}(2dn^{k^{\prime}})$, commuting gates compress groups to $\mathcal{O}(n^{k^{\prime}})$ ($k^{\prime}<k$)
VQNHE: $\mathcal{O}(2dn^{k})$, $n$ qubit, $\mathcal{O}(n^{k})$ operators, $d$ parameter
Commuting gates enable simultaneous measurements,
reducing cost to $\mathcal{O}(n^{k})$
Optimization
VQE: barren plateau in deep circuits
VQNHE: potential numerical instability
Neural-guided layer-wise optimization with gradual
amplitude transfer
Noise
VQE: affected by hardware noise and finite sampling noise
VQNHE: NN mitigates errors but amplifies sampling noise
Gradual amplitude transfer reduces sampling noise,
NN mitigates errors
Table 1: Key challenges in existing algorithms and sVQNHE’s mitigation strategies
Results
Theoretical Framework of sVQNHE
Figure 1: (1) The algorithm seamlessly integrates quantum circuits with neural networks. The quantum circuits interleaves diagonal layers ($W_{1}\cdots W_{n}$) with shallow, simultable layers ($G_{1}\cdots G_{n}$). Parameters of the $G$ layers are optimized classically using a gradual transfer mechanism, while those of the $W$ layers are adjusted via the parameter shift rule on the actual hardware. Leveraging the commuting properties of the diagonal layers, the number of measurements needed to compute gradients for the $d$ parameters of $W$ drops from $\mathcal{O}(2d)$ to $\mathcal{O}(1)$. (2) Illustration of sVQNHE’s advantages: the star represents the optimal solution, the green circle represents the initial solution space that is easy to explore by the NN, the orange region represents the solution space accessible to the PQC, and the black circle represents the coverage of the hybrid quantum-classical scheme. Among them, (2a) describes the suboptimal situation, that is, PQC cannot connect the classical region with the optimal solution, and the hybrid scheme is difficult to cover the optimal solution; (2b) shows the potential quantum advantage; (2c) describes the transition strategy from (2a) to (2b) through parameter adjustment. (3) Illustration of the energy optimization process using the hierarchical optimization scheme. When entering the next layer, the local optimal solution will be jumped out due to the newly introduced parameters.
We introduce sVQNHE, a hybrid quantum-classical variational algorithm designed to address key limitations of conventional VQAs, including high measurement cost, barren plateaus, limited expressive power with shallow circuits, and noise-induced errors. This framework builds on the original VQNHE approach by introducing a layer-wise, phase-amplitude decoupling strategy[24] (see Fig.1) that enhances trainability, scalability, and sampling robustness.
To elucidate the algorithm, we begin by summarizing the redesign of the quantum circuit and classical NN ansatzes. The overall (unnormalized) ansatz takes the form:
|$\ket{\psi}=F\prod_{i=0}^{L-1}W_{i}G_{i}\ket{0},$ (1)
where $F$ denotes a post-processing operator defined by a classical NN. The following paragraphs explain the design of the hybrid ansatz and the associated bidirectional optimization strategy. Finally, the algorithm is summarized in Alg. 1.
Quantum Circuit for Phase Structure Learning. The quantum circuit ansatz consists of alternating layers of parameterized gates:$\left(G_{i},W_{i}\right)$. $G_{i}$ are classically simulatable and non-diagonal layers (e.g., $R_{y}$ gates), primarily responsible for shaping the amplitude structure. $G_{1}$ is usually taken to be a layer of Hadamard gates to create a uniform superposition. $W_{i}$ are diagonal gates (e.g., $R_{z}$ and $R_{zz}$) designed to encode the phase structure of the wavefunction. These layers are added iteratively when greater expressivity is needed, and their parameters are optimized in a layer-wise fashion.
Classical Neural Network for Amplitude Modeling. The classical neural network is constrained to produce non-negative outputs, and is used to model the amplitude $F=\sum_{s}f(s)\ket{s}\bra{s}$ | of the quantum state in the computational basis. With the phase information delegated to the quantum circuit, the NN is focused on learning the probability distribution. Prior works have shown that classical neural networks are highly effective at this task, particularly for systems with smooth or structured amplitude landscapes [21, 24, 20].
Bidirectional Feedback Loop: Decoupling Amplitude and Phase. At each iteration $i$, a new block of paramterized quantum layers $(G_{i},W_{i})$ is added to the circuit. The process follows two key steps. (1) Forward Initialization Step: The classical NN operator $F$ from the previous iteration is approximated by the shallow quantum layer $G_{i}$, via a tractable classical optimization procedure as detailed in the Method section. This ensures that $G_{i}$ inherits the amplitude structure previously learned by the NN. (2) Hybrid Optimization: Once $G_{i}$ is initialized, the neural operator $F$ is reset and jointly optimized with the parameters of the new diagonal gate layer $W_{i}$, as done in the original VQNHE algorithm. This hybrid optimization simultaneously refines both the amplitude ($F$) and sign structure ($W_{i}$) in a tightly coupled quantum–classical feedback loop.
Algorithm 1 sVQNHE Hybrid Algorithm
1:$n$-qubit system, target Hamiltonian $H=\sum_{i}c_{i}P_{i}$
2:Optimized energy value $E_{\text{final}}$
3:Initialization
4:Initialize quantum state |$\ket{\psi_{0}}\leftarrow\ket{0}^{\otimes n}$
5:Initialize classical neural network parameters
6:Define diagonal operator $F_{1}\leftarrow\sum_{s}f(s)\ket{s}\bra{s}$ | $\triangleright$ $f(s)$: NN-predicted amplitude
7:Set $G_{1}\leftarrow H^{\otimes n}$triangleright Uniform superposition
8:Joint Optimization
[…]
10:    for $l\leftarrow 1$ to $L$ do $\triangleright$ Layer-wise optimization
11:       if $l>1$ then
12:          Update $\theta^{g_{l}}$ to minimize $\|G_{l}-F_{l-1}\|$triangleright Gradual Transfer Mechanism, see Method for detail
13:       end if
14:       $E_{l}\leftarrow 0$triangleright Reset layer energy
15:Hybrid Energy Estimation
16:       repeat
17:          Apply operator: |$\ket{\psi_{l}}\leftarrow W_{l}(\theta^{w_{l}})G_{l}(\theta^{g_{l}})\ket{\psi_{l-1}}$
18:          Construct hybrid state: |$\ket{\psi_{l,f}}\leftarrow F_{l}\ket{\psi_{l}}$
19:          for each Pauli term $P_{i}\in H$ do
20:             Compute $E_{P_{i}}\leftarrow\dfrac{\bra{\psi_{l,f}}P_{i}\ket{\psi_{l,f}}}{\innerproduct{\psi_{l,f}}{\psi_{l,f}}}$
21:             Accumulate energy: $E_{l}\leftarrow E_{l}+c_{i}\cdot E_{P_{i}}$
22:          end for
23:          Update NN parameters for $F_{l}$
24:          Update parameters: $\theta^{w_{l}}$triangleright Measurement Strategy, see Method for detail
25:       until $|\Delta E_{l}|<\epsilon_{\text{conv}}$triangleright Energy convergence
26:    end for
27:    Check global convergence $\triangleright$ Track $E_{l}$ across layers
28:end while
29:return $E_{L}$triangleright Final optimized energy
Adaptive Hybrid Ansatz with Controlled Expressivity
Conventional variational algorithms, such as the VQE and the QAOA, typically rely on predefined circuit ansatzes [37, 38]. These ansatzes, often inspired by physical principles like Trotterized dynamics in QAOA, impose a static architecture that can limit their effectiveness. For example, if the ansatz fails to capture the complex entanglement or sign structure of the target ground state, the optimization may converge to suboptimal local minima. To address this, adaptive methods like adapt-VQE [39, 10] iteratively construct the ansatz by selecting gates from a predefined pool. However, these approaches encounter significant hurdles: the gate pool may not include operations capable of generating sufficient entanglement, thus constraining expressiveness, and the iterative gate selection process requires a large number of quantum measurements, making it impractical for large-scale systems.
In contrast, VQNHE and sVQNHE employ a hybrid quantum-classical ansatz that effectively leverages both quantum and classical resources to represent a quantum state. This approach significantly expands the accessible quantum states with shallow quantum circuits. Nevertheless, the original VQNHE still assumes a fixed ansatz approach, which could clearly be optimized to further reduce circuit depth. To this end, sVQNHE adopts a neural-guided framework to systematically extend the circuit’s depth along with controlled expressiveness. Its shallow quantum layers $G_{l}$ are sequentially determined under the gradual guidance of a classical neural network.
A key feature of sVQNHE is its ability to achieve high expressiveness and stable convergence even with simple layers of $G_{i}$. For instance, take $G_{i}$ to be a layer of $R_{y}$ rotations on each qubit, can already enhannce the ansatz’s flexibility. This capability is rigorously supported by Theorem 1 (Expressiveness of the Quantum Circuit Ansatz), which analyzes the Lie algebra generated by specific sets of operators [40, 41, 12, 42, 43]. Consider the generator set $\mathcal{G}_{1}=\{Z_{D_{2j}}\mid j=1,2,\ldots,\lfloor\frac{m}{2}\rfloor\}\cup\{Y_{i}\mid i=1,2,\ldots,n\}$, where $Y_{i}$ are Pauli Y operators on individual qubits, and $Z_{D_{2j}}$ are Pauli Z operators on specific qubit subsets. The theorem proves that the corresponding Lie algebra is $\mathfrak{g_{1}}=\mathfrak{su}(2^{n-1})\oplus\mathfrak{su}(2^{n-1})$, a highly expressive structure with a dimension far exceeding that of alternative generator sets (e.g., $\mathcal{G}_{2}$). This large Lie algebra indicates that even layers primarily composed of $R_{y}$ gates, can explore a vast state space, enhancing the ansatz’s ability to approximate complex quantum states efficiently.
Moreover, the design of the quantum layers $G_{l}$ in sVQNHE is guided by structural constraints that ensure computational efficiency and scalability. Proposition 1 (Commuting Block Structure) provides a critical insight: if a block diagonal matrix $M=M_{1,1}\oplus\cdots\oplus M_{m,m}$ has distinct eigenvalues across its blocks, any matrix commuting with $M$ must preserve the same block structure. In the context of sVQNHE, this proposition informs the construction of $G_{l}$ layers. By respecting this commuting block structure, sVQNHE can increase circuit depth in a controlled manner, isolating commuting blocks (e.g., $W_{i}$).
Unlike rigid ansatzes (e.g., QAOA’s Trotterized structure) or gate-pool-dependent methods (e.g., adapt-VQE), sVQNHE avoids both architectural rigidity and measurement bottlenecks, achieves higher expressivity with less quantum resource demands.
Efficient Measurement and Trainability via Diagonal Gates and Neural Guidance
Traditional variational quantum algorithms face two intertwined challenges that limit scalability: high measurement costs and the emergence of barren plateaus. The first challenge arises from the need to estimate gradients by measuring non-commutative operators, a process requiring $\mathcal{O}(2d)$ measurements for $d$ trainable parameters. In large systems with $\text{poly}(n)$ operators, this results in a prohibitive $\mathcal{O}(2d\cdot\text{poly}(n))$ measurement overhead, which becomes computationally intractable as circuits scale. The second challenge, barren plateaus, manifests as exponentially vanishing gradients in high-dimensional parameter spaces, rendering optimization ineffective. These issues are exacerbated in deep circuits and systems with many qubits, where the parameter landscape grows increasingly complex and gradients vanish rapidly due to the curse of dimensionality.
The sVQNHE framework addresses these limitations through two synergistic innovations. First, it employs commuting diagonal gates, such as $R_{z}$ and $R_{zz}$ rotations, which enable simultaneous measurement of expectation values in one basis. This compresses the per-operator measurement cost to $\mathcal{O}(1)$, reducing the total overhead to $\mathcal{O}(\text{poly}(n))$ for $\text{poly}(n)$ operators (see details in the method section). This stands in stark contrast to the multiplicative $\mathcal{O}(2d\cdot\text{poly}(n))$ scaling of traditional VQNHE and conventional VQAs, allowing sVQNHE to tackle large-scale problems efficiently.
Secondly, our framework incorporates layer-wise neural guidance to decompose the optimization process into manageable subproblems, addressing a critical limitation of traditional layer-wise VQAs[44]. In conventional layer-wise VQAs, the absence of a global perspective on the optimization landscape renders sequential layer optimization prone to converging at local minima. In contrast, the sVQNHE framework employs a classical neural network to guide the optimization of each quantum layer, producing a hybrid state $|\psi_{l,f}\rangle$ that is more expressive than the state $|\psi_{l}\rangle$ obtained from hierarchical quantum circuits alone. This neural guidance not only helps avoid suboptimal local minima but also focuses parameter updates on one specific block each time, such as $W_{l}$ at the $l$-th iteration. This approach ensures that gradients remain measurable even in deep circuits, since the effective parameter space at each optimization step can always be restricted to a minuscule fraction of the full deep circuit. Together, these strategies circumvent the dual bottlenecks of measurement overhead and barren plateaus, establishing a scalable pathway for variational quantum algorithms in practical applications.
Sampling Noise Robustness
Under the finite sampling conditions, the mixture energy estimate in VQNHE requires normalization to counteract the non-unitary transformation induced by the neural network module $F$. The variance of the energy estimate $\hat{E}_{P_{i}}$ is governed by:
$\text{Var}(\hat{E}_{P_{i}})=\text{Var}\left(\bra{\psi_{L,f}}P_{i}\ket{\psi_{L,f}}\right)\cdot\text{Var}(g^{-1}(s,L)),$ (2)
where $g(s,L)=\frac{1}{n_{s}}\sum_{s_{i}}f^{2}(s_{i})$, and $s_{1},\ldots,s_{n_{s}}$ are i.i.d. samples drawn from the distribution $|\langle s|\psi_{L}\rangle|^{2}$. The expectation and variance of $g(s,L)$ are $\mathbb{E}[g(s,L)]=\mathbb{E}\left[f^{2}(s)\right]$ and $\text{Var}(g(s,L))=\frac{1}{n_{s}}\text{Var}\left[f^{2}(s)\right]$, where $s\sim|\langle s|\psi_{L}\rangle|^{2}$. This normalization process amplifies the sampling noise in the mixture energy estimate compared to the standard VQE, as quantified by $\text{Var}(g^{-1}(s,L))>1$. Notably, when the number of layer $L=1$ is fixed, this formulation form simplifies to that of the original VQNHE algorithm.
The sVQNHE algorithm introduces a layer-adaptive gradual transmission mechanism. As the quantum circuit depth $L$ increases, the algorithm shifts computational responsibility to the quantum component, progressively diminishing the weighting effect of $f(s)$across configurations $s$. This mechanism ensures:
$\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left(f^{2}(s)\right)\xrightarrow{L\uparrow}0.$ (3)
For sufficiently large $L$, $\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left(f^{2}(s)\right)<1$. By the uniform convergence of the series,
$\text{Var}(g^{-1}(s,L))=\sum_{i}\mu_{i}\text{Var}^{i}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left[f^{2}(s)\right],$ (4)
with $\mu_{i}>0$. This variance asymptotically approaches zero as $L\to\infty$. In our numerical experiment, as illustrated in Figures 3 and 4, the coefficient of variation of energy $\text{CV}\propto\sqrt{Var(g^{-1}(s,L))}$ decays with $L$, suggesting that the algorithm has better noise immunity under actual sampling constraints as $L$ increase. In summary, we eliminate a source of numerical instability (which subsequently affects the convergence rate) of the original VQNHE.
Experimental Results
To rigorously validate the accuracy and efficiency of our proposed algorithm for the ground-state preparation of a given Hamiltonian, we conducted systematic evaluations across four canonical manybody problems: the 1D J1-J2 model, the 2D Heisenberg model, the Transverse Field Ising Model (TFIM), and the Classical Ising model. These models were chosen to represent diverse physical systems with varying degrees of sign problem severity and ground-state complexity. We employed three methodologies, sVQNHE, VQE, and NNs, to benchmark their performance. Additionally, we integrated the encoding scheme proposed in Ref. [36] to demonstrate the superiority of sVQNHE over conventional VQAs (such as QAOA and VQE) in solving the MaxCut problem, a prototypical combinatorial optimization task.
Our experiments included a 6-qubit 1D J1-J2 model with $J1/J2=0.6$, a configuration known to exhibit a highly non-trivial ground-state phase structure [45]. This model serves as a representative case for systems with intricate sign signatures. We also tested a 9-qubit 2D Heisenberg model with periodic boundary conditions, which is known to incur severe sign-related challenges for numerical methods[46]. For comparison, we analyzed a 9-qubit 1D TFIM (solvable via Jordan-Wigner transform) and two 1D Ising models (6 and 12 qubits). Both TFIM and Ising models we chose are Sign-problem-free systems. Nevertheless, they still present non-trivial challenges in different ways. For example, they are relevant to the qubit-formulated combinatorial optimization problems. We evaluated our algorithm under four scenarios to assess its theoretical advantages and practical robustness:
  • Ideal Simulation: Quantified intrinsic accuracy, variability, and scalability under ideal conditions.
  • Real-World Simulation: Assessed robustness under realistic quantum conditions, combining depolarization-based gate errors (two-qubit: $0.005$; single-qubit: $0.0013$) with bitstring-sampling-based estimation of observables and gradients.
  • Scalability Under Finite Sampling: Demonstrated that sVQNHE’s advantage over existing methods[36] can be sustained with system size.
We emphasize that the results of all the so-called "non-ideal" experiments are based on the original sampling data (of bitstrings) from noisy quantum circuits, without implementing any quantum error mitigation procedures. As sVQNHE possesses an inherent and robust error mitigating capability [47] due to the fact that the hybrid quantum-neural ansatz can be directly tuned with the classical NN module during the noisy optimization process.
1. Ideal Simulation
We conducted a systematic evaluation of the performance of sVQNHE and other methods (VQNHE, VQE, NN, QAOA) focusing on accuracy, variability, and measurement efficiency. All the calculations were done in statevector-type exact numerical computations. For consistency, sVQNHE and VQNHE utilized the same non-negative real neural network component (MLP), while the NN method employed a real MLP with identical parameters and structure. The quantum components of sVQNHE and VQNHE were our single-layer sign ansatz and a hardware-efficient ansatz with a similar number of parameters, respectively. In Table 2, these are denoted as sign-NN and hea-NN, respectively. The VQE method used a hardware-efficient ansatz with varying layers, such as HEA2 for two layers. Similarly, the QAOA method was tested with different layers, for instance, QAOA10 for ten layers. All methods initialized quantum circuit parameters using independent uniform distributions $U(-2\pi,2\pi)$, optimized for the same number of steps, and averaged the estimated energy from 20 random initial points. The baseline algorithm in Table 2 is NN.
Simulating Frustrated Many-Body Systems
For the 1D J1-J2 model, which features a non-trivial ground state, the sVQNHE method (sign-NN) demonstrated significant improvements over the NN method, the VQE method (HEA2), and the VQNHE method (hea-NN). Specifically, Table 2 shows that after 2000 optimization steps, the sign-NN method reduced the mean absolute error (MAE) by $98.9\%$ and the variance by $99.6\%$ for the 6-qubit J1-J2 model compared to the NN method. Additionally, Figure 2 illustrates the scaling results (from 6 qubits to 12 qubits) of sign-NN and NN with the same number of optimization steps to achieve the minimum error. The advantage of sign-NN increases with the number of qubits. In contrast, while hea-NN also reduced the MAE by $75.1\%$ compared to NN, its variance was larger than that of NN (variance: hea-NN > NN > sign-NN). Furthermore, due to the non-commutative structure of hea, the measurement amount for sign-NN was significantly smaller than that for hea-NN under similar parameter amounts. These three points—improved accuracy, reduced variance, and reduced measurement amount—demonstrate the superiority of sVQNHE over the original VQNHE algorithm. Regarding convergence speed, sign-NN significantly outperformed HEA2. To achieve the same error threshold, HEA2 required 12.6 times more optimization steps on average than sign-NN. In the most efficient cases, HEA2 needed 1676 steps to reach the same accuracy, while sign-NN only required 89 steps. Additionally, the probability of the sign-NN successfully reaching the target accuracy was 12 times higher than that of HEA2, highlighting its excellent reliability and efficiency (for details, please refer to the supplementary material). Similar results were observed in the Heisenberg model, further verifying the capability of sVQNHE to handle systems affected by the sign problem.
Model Method $R_{MAE}$ $R_{Var}$
6-qubit 1D J1-J2 sign-NN -0.989 -0.996
hea-NN -0.751 1.955
HEA2 -0.916 -0.132
9-qubit 2D Heisenberg HEA1 -0.331 0.554
hea-NN -0.119 33.793
sign-NN -0.872 -0.832
9-qubit 1D TFIM HEA1 -0.290 0.019
sign-NN -0.640 -0.728
9-qubit 1D Ising HEA1 0.129 0.909
QAOA5 20.674 15.583
QAOA10 1.205 11.835
sign-NN -0.416 -0.455
12-qubit 1D Ising HEA1 6.822 10.002
QAOA10 77.012 80.404
sign-NN -0.811 -0.979
Table 2: Performance Analysis of sVQNHE Algorithm. Here, the range of $R_{MAE}$ and $R_{Var}$ is $(-1,+\infty)$. The closer to $-1$, the better (see Equations (12) and (13) in the Method section for details). The baseline algorithm NN model is a real Multi-Layer Perceptron (MLP). HEA1 and HEA2 represent VQE methods employing 1-layer and 2-layer Hardware-Efficient Ansatz, respectively. QAOA5 and QAOA10 represent the QAOA method with layers 5 and 10, respectively. The Sign Ansatz employs $R_{zz}$ gates to connect all edges of the corresponding physical model.
Figure 2: Scaling test for the 1D J1-J2 model. The x-axis shows the number of qubits, while the y-axis displays the average relative error from the sVQNHE method (sign-NN, blue line) and the NN method (NN, orange line) after 20 runs of 2000 steps each. The green line illustrates the difference in average relative error between the NN and sign-NN methods as the qubit count increases. Shaded regions represent the relative error bar scaled by 0.5.
Simulating Sign-Problem-Free Many-Body Systems
Although the above results highlight the advantages of sVQNHE in sign-frustrated quantum models, the same architectural design is effective even in sign-free systems, where phase modeling is trivial for the ground state itself, but the ability to explore a broader Hilbert space (as defined by the PQC-generated phase strcutures) turns out to be beneficial for the optimization convergence in the numerical investigated presented in this subsection. More precisely, please see the last three rows of Table 2.
In our numerical experiments on the Ising model, the sign-NN approach demonstrates substantial improvements over the NN in both error reduction and variance compression. At 9 qubits, sign-NN achieves a $41.6\%$ reduction in MAE compared to NN, indicating a significant enhancement in accuracy. This advantage becomes even more pronounced at 12 qubits, where sign-NN reduces MAE by $81.1\%$ relative to NN, showcasing its superior scalability with increasing system size. Similarly, in terms of variance, sign-NN outperforms NN by reducing variance by $45.5\%$ at 9 qubits. This gap widens dramatically at 12 qubits, with sign-NN achieving a $97.9\%$ reduction in variance compared to NN, highlighting its ability to deliver consistent and reliable results as the system scales. When benchmarked against the QAOA in the 9-qubit Ising model, sign-NN maintains its performance edge. A 5-layer QAOA configuration requires a comparable number of per-iteration measurements to sign-NN but yields a higher MAE, indicating lower accuracy. Extending QAOA to 10 layers improves its accuracy, yet it still falls short of sign-NN’s performance. This trend persists and becomes more evident at 12 qubits, where sign-NN’s scalability further amplifies its advantage over QAOA. The consistency of these results extends to the TFIM. The superior performance of sign-NN can be attributed to its carefully designed ansatz, optimized with neural network guidance. Unlike QAOA, which uses $R_{x}$ gates and constrains all $Z$- and $X$-related gates within a layer to share a single parameter, the Sign Ansatz employs $R_{zz}$ and $R_{y}$ gates with independent parameters. This design, as supported by Theorem 1, generates a higher-dimensional Lie algebra ($\mathfrak{g}_{1}$) compared to QAOA’s ($\mathfrak{g}_{2}$), enabling sign-NN to explore a broader unitary space through multi-qubit $Z$-phase adjustments. The phase-focusing architecture of the Sign Ansatz, which leverages multi-qubit $Z$-gates to refine phase information, allows for efficient navigation of high-dimensional solution spaces. When combined with neural network optimization, this approach achieves high accuracy and low variance at a reduced computational cost, making sVQNHE particularly well-suited for large-scale quantum optimization problems.
2. Real-World Simulation
Next, we consider how robustly sVQNHE performs when there exists both gate errors (for e.g. depolarization channels) and uncertainty induced by the finite sampling of bitstrings instead of statevector calculations as in the previous section. We evaluated sVQNHE against VQE, adopting two distinct optimization strategies, layered-VQE (layerwise optimization) and standard-VQE (joint optimization of all parameters), using an identical ansatz circuit. sVQNHE employed non-negative real neural networks, while both VQE methods utilized the same Sign Ansatz without NN augmentation. For all VQE schemes, the resource consumption was saved by measurement based on commuting groups. All simulations capped iterations at 400.
Figure 3: Performance evaluation of sVQNHE and VQE variants under a 3-qubit 1D Heisenberg Hamiltonian. Energy (y-axis) versus iteration count (x-axis), with the ground state energy marked by a black dashed line. For sVQNHE and layered-VQE (layered optimization), blue and orange lines correspond to the first and second evolutionary layers, respectively, where each layer introduces new variational parameters. For standard-VQE (overall optimization), the lines reflect energy progression across sequential steps under fixed parameterization. For sVQNHE, coefficients of variation after stabilization (final 50 iterations per layer) are $cv_{1}=0.01618$ and $cv_{2}=0.015912$, indicating a reduction in variability in the second layer.
Figure 4: Energy convergence of sVQNHE and VQE variants under a noisy 3-qubit 1D Heisenberg Hamiltonian. Iteration count (x-axis) versus energy (y-axis). The solid black dashed line marks the ground state energy value. For layered optimization (sVQNHE, layered-VQE), blue and orange lines denote first- and second-layer parameter updates, respectively; standard-VQE (overall optimization) reflects sequential parameter refinement. For sVQNHE, coefficients of variation after stabilization (final 50 iterations per layer) are $cv_{1}=0.01534$ and $cv_{2}=0.01487$, indicating a reduction in variability in the second layer.
As shown in Figures 3 and 4, sVQNHE consistently achieved faster convergence than VQE, both with and without device noise. While layered optimization inherently reduces computational resource demands, mitigates noise sensitivity via shallower circuits, and often accelerates convergence, it is not applicable to the general VQE framework due to premature convergence to local optimal solutions. Specifically, the 1-layer Sign Ansatz fails as a viable initial parameterization for VQE’s layered optimization, as evidenced by the divergent performance of layer-VQE and standard VQE in our tests. However, integrating NNs into the Sign Ansatz (as in sVQNHE) circumvents this limitation, enabling effective layered optimization by delegating partial modeling duty to the neural network. Post-convergence analysis of sVQNHE revealed that the coefficient of variation (CV) for the second optimization layer remained lower than that of the first layer across all noise conditions. This validates the efficacy of our gradual transfer mechanism, which systematically refines solutions while maintaining stability under noise.
3. Scaling Under Sampling Errors
Finally, we investigate whether the observed advantage of sVQNHE can be sustained as the system size increases under the real-world setting of extracting PQC state by sampling bitstrings on quantum circuits. An example is provided in Figure 5. We evaluate the solution quality and measurement cost of the MaxCut problem on Erdős–Rényi random graphs with an edge connection probability of $0.3$ as the graph size increases. To establish this advantage, we compare sVQNHE against the Sign Ansatz VQE (sign-VQE) and the recently proposed brickwork-VQE [36], all evaluated under the same polynomial encoding model. This model encodes the $m$ binary variables of the MaxCut problem into the Pauli correlations across $k$ qubits, enabling the representation of $m=O(n^{k})$ variables with only $n$ qubits, where $k>1$ is an integer. Unlike the direct encoding schemes like QAOA, which require $m$ qubits—one per variable—this approach achieves polynomial compression in the qubit count. The desired solution of the combinatorial problem is now encoded in terms of a non-classical Hamiltonian. The compression ratio depends on the polynomial order $k$, allowing us to tackle larger MaxCut instances with fewer qubits. All methods are evaluated under a unified encoding framework with $k=2$ to ensure fairness.
To evaluate solution quality across varying scales, we conducted experiments with 1,200 measurement shots, averaging results over multiple trials. As shown in Figure 5(Left), scaling sVQNHE parameters—through enhancements in the NN component or other optimizations—consistently widens its performance gap over competing methods. Specifically, sVQNHE, which employs NN-guided hierarchical optimization and NN post-processing, significantly outperforms sign-VQE and brickwork-VQE. Sign-VQE adopts the final PQC structure from sVQNHE (a two-layer sign ansatz with two diagonal and two non-diagonal layers) but omits the NN-guided optimization and post-processing used in sVQNHE. Brickwork-VQE, in contrast, utilizes a VQE with a brickwork ansatz, as detailed in [36]. Figure 5(Left) reveals that, without NN guidance and post-processing (sign-VQE vs. brickwork-VQE), the sign ansatz shows no clear advantage over the brickwork ansatz, with comparable solution quality and similar measurement costs. However, sVQNHE’s hybrid approach - integrating hierarchical optimization guided by NN and NN post-processing to Sign Ansatz - yields a marked improvement. For instance, the performance gap between sVQNHE and baseline methods increases by approximately $6\%\approx(0.96-0.90)/0.90$ for 9-vertex graphs and $19\%\approx(1-0.84)/0.84$ for 45-vertex graphs when using parameter set 2. Additionally, Figure 5(Right) shows that sVQNHE significantly reduces single-step measurement consumption compared to both sign-VQE and brickwork-VQE. With parameter set 2, it achieves $86\%\approx(72-10)/72$ and $85\%\approx(306-46)/306$ improvements in quantum resource efficiency for 9- and 45-vertex systems, respectively. Additionally, in the Supplementary Information, we provide a preliminary comparison of resource consumption for larger instances, further highlighting sVQNHE’s advantages in scalability and efficiency under large-scale conditions. This underscores its potential for large-scale quantum applications.
Figure 5: Scaling test of sVQNHE using the MaxCut problem as an example. (Left) The relative solution quality for the MaxCut problem, comparing parameter sets 1 and 2 across different problem sizes. The methods sign-VQNHE, brickwork-VQE, and sign-VQE correspond to the sVQNHE method with a 2-layer sign ansatz, the VQE method with brickwork ansatz, and the VQE method with 2-layer sign ansatz, respectively. The depth of brickwork ansatz is determined according to the reference [36], that is, the circuit depth is sublinearly related to the number of vertices in the graph. For sign-VQE, the ansatz is the same as the circuit finally generated by the corresponding sVQNHE scheme finally generated, utilizing a two-layer sign ansatz with two diagonal and two non-diagonal layers. The value range of relative solution quality is $[0,1]$. The closer to 1, the better the performance is compared with other algorithms (see Equation (14) in the Method section for details). For example, in the case of 45 vertices, the quality of the solution of sVQNHE is relatively optimal regardless of whether the parameter set is 1 or 2. (Right) The single-step measurement consumption for the same problem sizes. Parameter set 2 features an increased number of parameters in the NN component and doubles the number of iterations compared to parameter set 1. The circuit depth in the brickwork architecture scales sub-linearly with the number of vertices.
Discussion
We present sVQNHE, a hybrid quantum-classical framework that strategically optimizes the roles of quantum and classical resources. Compared to the original VQNHE, our approach significantly improves measurement efficiency, substantially suppresses the numerical instability of the neural network, and enhances overall convergence.
Notably, sVQNHE leverages commuting diagonal gates to encode quantum sign structures, allowing efficient gradient estimation with far fewer measurements. The classical neural network is tasked with learning wavefunction amplitudes and guiding parameter updates, effectively sidestepping the barren plateau issue commonly seen in deep quantum circuits. This division of labor enables a layer-wise optimization strategy that remains stable even under noise and sampling limitations.
Our numerical simulations demonstrate compelling improvements across a broad class of computational problems. In the 6-qubit J1-J2 model, the sVQNHE method with a single-layer assumption achieves more than 7-fold the average accuracy and 19-fold the convergence speed of the 2-layer hardware-efficient VQE. Compared with the baseline NN, the average accuracy is improved by nearly two orders of magnitude. In sign-free models such as the TFIM and Ising chains, sVQNHE consistently outperforms QAOA and VQE, achieving broader expressivity and better scalability. For combinatorial optimization, such as MaxCut on 45-node graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$ compared to brickwork ansatz-based VQE.
Taken together, these results highlight sVQNHE’s broad applicability, efficiency, and scalability. By tightly integrating neural networks with quantum circuits along with a revised role that each component plays in the new algorithm, sVQNHE offers a practical path toward quantum utility on NISQ and early fault-tolerant devices. Looking ahead, this framework can be further extended by incorporating advanced machine learning models and better measurement strategies, and applied to a wider class of many-body simulation and large-scale optimization problems. Extending sVQNHE to tackle PQC-based quantum machine learning is also conceptually straightforward. As quantum hardware continues to evolve, we expect sVQNHE to remain a strong candidate for bridging theoretical capabilities with real-world impact.
Methods
The Hamiltonian of J1-J2 Model and Heisenberg Model
For the 2D Heisenberg models are constructed with non-periodic boundary conditions, whose Hamiltonian can be written as,
$H=h_{i}\sum_{i}Z^{i}+J_{ij}\sum_{edges}X^{i}X^{j}+Y^{i}Y^{j}+Z^{i}Z^{j}.$ (5)
In our numerical simulation, we use the coefficient setting of $(h_{i},J_{i,j})$ to be $(1,0.4)$ in $9$-qubit 2D Heisenberg model and $(1,1)$ in $3$-qubit Heisenberg model.
For the Hamiltonian of the n-qubit 1D J1-J2 spin model with non-periodic boundary conditions, it can be written as
$H=\sum_{i=1}^{n-1}J_{1}(X^{i}X^{i+1}+Y^{i}Y^{i+1}+Z^{i}Z^{i+1})+\sum_{i=1}^{n-2}J_{2}(X^{i}X^{i+2}+Y^{i}Y^{i+2}+Z^{i}Z^{i+2}).$ (6)
In our numerical simulation, we use the coefficients $J_{1}=1$ and $J_{2}=0.6$, since for the $J_{2}>0.5$ the sign problem is non-trivial for the J1-J2 model.
Detail of the Theoretical Framework
Theorem 1
(Expressiveness of the Quantum Circuit Ansatz)[40, 41, 12, 42, 43] Let $\mathfrak{g_{1}}$ and $\mathfrak{g_{2}}$ be the Lie algebras corresponding to the generators $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, where:
$\mathcal{G}_{1}=\{Z_{D_{2j}}\mid j=1,2,\ldots,\lfloor\frac{m}{2}\rfloor\}\bigcup\{Y_{i}\mid i=1,2,\ldots,n\},$
$\mathcal{G}_{2}=\left\{\sum Z_{D_{2j}}\mid j=1,2,\ldots,\lfloor\frac{m}{2}\rfloor\right\}\bigcup\left\{\sum X_{i}\mid i=1,2,\ldots,n\right\},$
where $m\in\{2,3,\ldots,n\}$, $X_{i}$ and $Y_{i}$ are the Pauli strings with Pauli X and Pauli Y operators at the ith position, respectively, $D_{2j}$ is the set of all 2j-combinations from the set $\{1,2,\ldots,n\}$, and $Z_{D_{2j}}$ is the Pauli string with the Pauli Z operator on the set of positions $D_{2j}$. Then, the following statements hold:
  • $\mathfrak{g_{1}}=\mathfrak{su}(2^{n-1})\oplus\mathfrak{su}(2^{n-1})$.
  • $dim(\mathfrak{g_{2}})<dim(\mathfrak{g_{1}})$.
Proposition 1
(Commuting Block Structure) Let $M$ be a block diagonal matrix $M=M_{1,1}\oplus\cdots\oplus M_{m,m}$ with distinct eigenvalues across blocks. Then, any matrix commuting with $M$ must share its block structure.
Gradual transfer mechanism
Our algorithm enforces strict classical-to-quantum information transfer by iteratively aligning the classically simulable $G_{l}$ (shallow quantum circuit blocks) with the classical part $F_{l-1}$. This can be performed by different schemes, such as:
1. Frobenius norm minimization (global): Minimize the full matrix distance: $|G_{l}-F_{l-1}|_{F}^{2}$.
2. Test state projection consistency (local): For a set of Haar-random test states $\{|\psi_{\text{test}}^{k}\rangle\}$, enforce consistency in the projected probability distribution
$P_{G}^{k}(s)=|\langle s|G_{l}|\psi_{\text{test}}^{k}\rangle|^{2},\quad P_{F}^{k}(s)=|\langle s|F_{l-1}|\psi_{\text{test}}^{k}\rangle|^{2}$ (7)
by minimize expected KL divergence $\mathbb{E}_{k}\left[\text{KL}(P_{G}^{k}\|P_{F}^{k})\right]\to 0$. This test state sampling provides a weaker but practical consistency condition, avoiding direct optimization of $G_{l}$.
Measurement Strategy and Cost
In our layered optimization strategy, we focus on performing measurements for any given layer. To estimate the measurement cost of the Sign Ansatz, we use the parameter shift rule mentioned in Equation (8) to compute the gradient of each parameter in the circuit. The parameter shift rule is defined as follows:
$\gradient_{\theta}f(x;\theta)=\frac{1}{2}[f(x;\theta+\frac{\pi}{2})-f(x;\theta-\frac{\pi}{2})].$ (8)
To maximize the efficiency of executing the parameter shift rule, we exploit the commutative property of $W(\theta^{W})$. During iterative parameter updates, this approach simplifies the process: we only need to add the corresponding quantum gate after circuit W for each updated parameter. Specifically, for the parameters in $W(\theta^{W})=W_{1}(\theta_{1}^{W})\cdots W_{n_{W}}(\theta_{n_{W}}^{W})$, we have
$\frac{\partial E_{P_{i}}(\theta^{W})}{\partial\theta_{k}^{W}}=\frac{1}{2}(E_{P_{i}}(\theta_{k}^{W}+\frac{\pi}{2})-E_{P_{i}}(\theta_{k}^{W}-\frac{\pi}{2})).$ (9)
Since $W_{k}$ and $F$ are all commutative, and $E_{P_{i}}(\theta^{W})=\bra{\psi_{0}}W^{\dagger}(\theta^{W})F^{\dagger}P_{i}FW(\theta^{W})\ket{\psi_{0}}$, we have
$E_{P_{i}}(\theta_{k}^{W}\pm\frac{\pi}{2})$
$=\bra{\psi_{0}}W^{\dagger}(\theta^{W})F^{\dagger}W_{k}^{\dagger}(\pm\frac{\pi}{2})P_{i}W_{k}(\pm\frac{\pi}{2})FW(\theta^{W})\ket{\psi_{0}}.$ (10)
Equation (9) and (Measurement Strategy and Cost) imply that to obtain the gradient of $\theta^{W}$, we only need to add the measurement of Puali string $L_{i}=W_{k}^{\dagger}(\frac{\pi}{2})P_{i}W_{k}(\frac{\pi}{2})$.
Note that any measurement of $Z\in L_{i}$ can be done by adding a factor $\prod_{i\in ind_{z}}(1-s_{i})$ to the coefficient $(1-2s_{0})$, where $ind_{z}$ is the position of Z in $L_{i}$ and $s_{i}$ is the Boolean value of the i index, so only changes in X and Y in $L_{i}$ will cause the number of measurements to increase. Let $P_{i}$ be a Pauli string containing exactly $|p_{1}|$ Pauli X or Y operators. Let $W_{k}$ be the rotation gate corresponding to $Z_{p}$ with $|p_{2}|$ operators. Since the transformations between X and Y can occur between $L_{i}$ and $P_{i}$ only if $|p_{1}\cap p_{2}|$ is odd or $|p_{1}\cap p_{2}|=|p_{2}|$, then for a Hamiltonian with $|p_{1}|\leq 2$, such as the 1D J1-J2 model, the total amount of measurement we need is $O(m_{H})$, where $m_{H}$ is the number of Pauli strings $P_{i}$ in $H$. Here, $p_{1}$ and $p_{2}$ are non-empty subsets of the power set of all positions of the n qubit system.
In summary, due to the commutation properties of the Sign Ansatz, our algorithm requires fewer measurements compared to general algorithms that rely on parameterized quantum circuits. For instance, considering the 1D J1-J2 model presented below (11), if $|p|\geq 2$ in the Sign Ansatz, the total number of measurements in a single iteration of our algorithm is equal to $13(n_{1}+n_{2})+$ 1, where $n_{1}$ is the number of $\langle i,j\rangle$, and $n_{2}$ is the number of $\langle\langle i,j\rangle\rangle$.
$H=J_{1}\sum_{\langle i,j\rangle}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{j}^{z})+J_{2}\sum_{\langle\langle i,j\rangle\rangle}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y}+\Delta_{2}\sigma_{i}^{z}\sigma_{j}^{z})+B_{H}\sum_{i}\sigma_{i}^{z}.$ (11)
Here $\sigma^{x,y,z}$ are the Pauli matrices, $\Delta_{1}$ and $\Delta_{2}$ are the anisotropy parameters for nearest-neighbor and next-nearest-neighbor interactions, and $B_{H}$ is the strength of an external magnetic field. $\langle i,j\rangle$ and $\langle\langle i,j\rangle\rangle$ denote sums over nearest-neighbor and next-nearest-neighbor spin pairs, respectively.
The details of our numerical simulation
We use TensorCircuit[48] to efficiently obtain all the quantum circuits and the neural network numerical results with the following Ansatz setting, and the details of the VQNHE algorithm are in the TensorCircuit Demo.
[…]
The Sign Ansatz design equivalents for the sampling and noise versions of the numerical results with finite shots are detailed as follows. A combination of $YM$, $YP$, $C_{z}$, and $R_{z}$ gates equivalently replaces the $R_{zz}$ gate. Similarly, the $R_{y}$ gate is equivalently replaced by a combination of $XP$, $XM$, and $R_{z}$ gates. Here, $YM$ and $YP$ denote positive and negative rotations by $\frac{\pi}{4}$ around the Y-axis of the Bloch sphere, respectively, while $XM$ and $XP$ denote positive and negative rotations by $\frac{\pi}{4}$ around the X-axis of the Bloch sphere, respectively. The two-qubit gates used in our cases connect the edges of the model (1D J1-J2 model and 2D Heisenberg model).
[…]
In the design of Hardware Efficient Ansatz, where the $CNOT$ is only connecting the two nearest qubits (2D Heisenberg Model), n is the number of repetitions. Note that after repeating the above block n times, one more layer of $R_{y}$ and one more layer of $R_{z}$ are added. For example, for HEA2 used in this paper, there are three $R_{y}$ layers, three $R_{z}$ layers, and two $CNOT$ layers. Similarly, the $CNOT$ gate in the numerical simulation of the finite-shot version is equivalently replaced by a combination of $YM$, $YP$, and $C_{z}$ gates.
Metrics
In the experiment, we use the following metrics to evaluate the performance of the method:
  • Mean absolute error ratio $R_{MAE}$:
    $R_{MAE}=\frac{MAE_{1}-MAE_{0}}{MAE_{0}}.$ (12)
    The metric $R_{MAE}$ represents the relative reduction in mean absolute error (MAE) of different algorithms compared to the baseline algorithm (NN) with the same number of iterations. Here $MAE_{0}$ represents the MAE of the baseline algorithm (NN) being compared. The smaller the $R_{MAE}$ value, the better the compared algorithm is relative to NN.
  • Variance ratio $R_{Var}$:
    $R_{Var}=\frac{Var_{1}-Var_{0}}{Var_{0}}.$ (13)
    The metric $R_{Var}$ represents the relative reduction in the variance of different algorithms compared to the baseline algorithm (NN) given the same number of iterations. Among them, $Var_{0}$ represents the variance of the baseline algorithm (NN) being compared. The smaller the $R_{Var}$ value, the more stable the compared algorithm is relative to NN, and the smaller the variability of the results.
  • Estimated approximation ratio $R_{e}$ for the MaxCut problem:
    $R_{e}=\frac{MaxCut_{algo}}{MaxCut_{max}}.$ (14)
    The metric $R_{e}$ provides an estimated approximation ratio based on the best-known result for the MaxCut problem, which measures the relative solution quality. Here, $MaxCut_{max}$ represents the maximum MaxCut value achieved across all evaluated algorithms, while $MaxCut_{algo}$ denotes the MaxCut value obtained by the specific algorithm under assessment. A value of $R_{e}$ approaching $100\%$ indicates superior performance relative to the other algorithms compared.
Acknowledgements
CYH acknowledges support from National Ke Research and Development Program of China (2024FA1306400), Natural Science Foundation of China (22373085). AH gratefully acknowledges the sponsorship from Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. CityU 11200120), City University of Hong Kong (Project No. 7005615, 7006103), CityU Seed Fund in Microelectronics (Project No. 9229135), and Hon Hai Research Institute (Project No. 9231594). YYS acknowledges support from the National Natural Science Foundation of China (Grant No. 12404575), the Shandong Provincial Natural Science Foundation (Grant No. ZR2022LLZ008), the Natural Science Foundation of Shanghai (Grant No. 23ZR1469600), and the Shanghai Sailing Program (Grant No. 23YF1452600).
Code and data availability
The source code and data associated with this project are publicly available. They can be accessed at the following GitHub repository: https://github.com/renmengzhen/sVQNHE
Supplementary
Optimizing Quantum Resources with Sign Ansatz VQNHE
In this section, we evaluate the quantum resource efficiency of the sVQNHE framework against the Hardware Efficient Ansatz VQE. First, we test different ansatz on the 9-qubit 2D Heisenberg model with the initial state $\ket{++...++}$. The results are shown in Figure 6, it can be seen that sVQNHE uses very few quantum resources to converge to the desired accuracy compared to both neural networks and VQE, which implies that the sVQNHE reaches the desired accuracy with fewer quantum resources compared to VQE in the number of circuits, and the number of parameters. Besides, since the zero quantum resource results indicate that without the assistance of Sign Ansatz the neural network itself cannot reach good accuracy, we can secondly claim that this framework can greatly improve the expressiveness of the neural network with limited quantum resources. In addition, we also verified the quantum resource efficiency of sVQNHE relative to VQE on the J1-J2 model. The results are shown in the Table 3.
Figure 6: 9-qubit Heisenberg Model Result. The x-axis shows sorted results by the converged energy with different random initial parameters. The y-axis is the energy difference from the ground state. The red, green, and blue dashed lines are responding to Hardware Efficient Ansatz with repetitions equal to 1, 2, and 3, and the solid line is Complex MLP(3,3) with Sign Ansatz, where the total number of parameters in the neural network is 317. Both of them start from the initial state $\ket{+...+}$. The dotted line $99.99\%$ Energy is 0.0001 times the ground state energy.
Table 3: Convergence Efficiency Comparison for the 6 qubit J1-J2 Model
Method Average Optimization Steps Success Probability (%) Summary
sVQNHE 132.67 60 faster convergence with higher reliability
VQE (2-layer HEA) 1676.00 5 more steps and has a lower success rate
Explanation: This table compares the convergence efficiency of sVQNHE and VQE (with a 2-layer hardware-efficient ansatz) for the 6-qubit J1-J2 model. The metrics include the average number of optimization steps required to reach 99.45% of the ground state energy and the probability of success. sVQNHE outperforms VQE by requiring fewer steps (132.67 vs. 1676) and achieving a higher success rate (60% vs. 5%), demonstrating superior efficiency in quantum resource utilization.
Sampling Efficiency Verification
This section presents a compelling argument for performing sampling from the quantum circuit component rather than the classical neural network component that describes the amplitude. Our objective is to verify the computational advantages in the main text argument. As illustrated in Figure 7, we compare the convergence curve of our model with that shown in Figures 3 and 4 of the main text. The key difference is that the set of bitstrings $s$ here is obtained by sampling from the classical neural network component. Compared to the results in Figures 3 and 4 of the main text, we observe that performing sampling from the quantum circuit component is better; fewer shots can yield more stable iterations and faster convergence. This suggests that our quantum circuit can effectively sample from complex probability distributions that are intractable for classical computers, thereby providing a high-quality solution space that closer approximates the true solution space.
Figure 7: Sampling Efficiency Verification under 2D Heisenberg Hamiltonian. The figure illustrates the results obtained under varying numbers of qubits and shots. Sampling is conducted from the classical neural network component. The x-axis denotes the number of iterations, while the y-axis represents the energy. The solid dotted line indicates the ground state energy value of the corresponding model.
Preliminary Resource Consumption Analysis for Larger Instances
Table 4: Resource Comparison for Larger Instances
Qubits Encoding Vertices
brickwork-VQE
measurement cost
Comparison
sVQNHE
measurement cost
17 2 408 3000 $>$ 409
3 2040 24150 $>$ 4081
30 2 1305 9612 $>$ 1306
3 12180 141246 $>$ 24361
Explanation: This table compares the measurement costs of sVQNHE and brickwork-VQE under 2nd and 3rd order polynomial encodings for 17 and 30 qubits. The vertex count represents the maximum number of graph vertices for the maxcut problem corresponding to each qubit count and encoding scheme. Measurement cost reflects the number of measurement circuits required per iteration. sVQNHE consistently requires fewer circuits, demonstrating superior efficiency in quantum resource utilization.
In this section, we compare the measurement requirements for a single iteration of the brickwork-VQE and sVQNHE schemes under 2nd and 3rd order polynomial encodings[36], with equal shots per operator. As shown in the Table. 4, the savings in measurements, defined as the reduction when using sVQNHE compared to brickwork-VQE, scale linearly with the number of vertices in the graph. While this scaling is linear, the absolute savings become significantly large for graphs with many vertices, resulting in a substantial efficiency advantage for sVQNHE in large-scale quantum systems.

### v2

@@
 Abstract
-Variational quantum algorithms hold great promise for unlocking the power of near-term quantum processors, yet they are frequently hindered by high measurement costs, barren plateaus, and challenging optimization landscapes.
-Here, we introduce sVQNHE, a neural-guided variational quantum algorithm that decouples amplitude and sign learning across classical and quantum modules, respectively.
-Our approach employs shallow quantum circuits composed of commuting diagonal gates to efficiently model quantum phase information, while a classical neural network learns the amplitude distribution and guides circuit optimization in a bidirectional feedback loop.
-This hybrid quantum-classical synergy not only reduces measurement costs but also achieves high expressivity with limited quantum resources and improves the convergence rate of the variational optimization.
-We demonstrate the advancements brought by sVQNHE through extensive numerical experiments (including realistic noisy simulations and sampling-based simulations).
-For the 6-qubit J1-J2 model — a prototypical system with a severe sign problem for Monte Carlo-based methods — it reduces the mean absolute error by $98.9\%$ and suppresses variance by $99.6\%$ relative to a baseline neural network, while requiring nearly $19x$ fewer optimization steps than a standard hardware-efficient VQE.
-Furthermore, for MaxCut problems on 45-vertex Erdős–Rényi graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$.
-Importantly, this framework is designed to be scalable and robust against hardware noise and finite-sampling uncertainty, making it well-suited for both current NISQ processors and future high-quality quantum computers.
-Our results highlight a promising path forward for efficiently tackling complex many-body and combinatorial optimization problems by fully exploiting the synergy between classical and quantum resources in the NISQ era and beyond.
+Variational quantum algorithms are a leading candidate for near-term quantum computing, yet their scalability is hindered by high measurement costs, barren plateaus, and noise sensitivity.
+These challenges stem largely from requiring quantum circuits to learn both probability amplitudes and phase (sign) structure simultaneously—a demanding regime that increases circuit depth and sampling overhead.
+Here we introduce sVQNHE, a hybrid quantum–classical framework that explicitly separates these roles: a classical neural network models the amplitude distribution, while shallow, commuting diagonal quantum circuits learn only the phase structure.
+Furthermore, sVQNHE employs an iterative amplitude-transfer mechanism that progressively integrates the classical distribution into the quantum ansatz—while leveraging commuting diagonal gates to reduce gradient measurement cost to $\mathcal{O}(1)$ per operator—ensuring the neural network models only a residual correction on a highly trainable energy landscape with polynomially bounded gradient variance.
+Across frustrated quantum spin models, molecular Hamiltonians, and combinatorial optimization benchmarks, sVQNHE significantly outperforms conventional variational quantum eigensolvers with comparable parameter counts, converging faster and demonstrating improved robustness under finite sampling and noise.
+Notably, using qubit-efficient encodings, sVQNHE matches the state-of-the-art classical Free-Energy Machine on a 1,485-vertex MaxCut instance using Pauli Correlation Encoding, and substantially outperforms all baselines on Maximum Clique problems with 135 vertices.
+These results establish that explicit amplitude–phase separation, with quantum resources dedicated mainly to learning phase structure, provides a principled and scalable design strategy for hybrid quantum–classical algorithms under near-term noisy hardware constraints.
 Introduction
-Quantum computing promises transformative advances in a variety of applications, from combinatorial optimization to ground-state preparation and dynamical simulation of many-body systems [1, 2].
-However, large-scale quantum circuits remain challenging to implement due to hardware limitations, gate errors, and limited qubit coherence [3].
-One of the most urgent goals in the noisy intermediate-scale quantum (NISQ) era, and the subsequent early fault-tolerant era [4] (Early FTQC), is to maximize computational utility within these constraints by employing shallow circuits and error-mitigated designs [5].
-Among the many strategies proposed to enable practical applications under these conditions, hybrid quantum–classical algorithms, particularly variational quantum algorithms (VQAs) [6, 7, 8], stand out.
-VQAs combine parameterized quantum circuits (PQC) with classical optimization routines, using quantum circuits as ansatzes to solve problems that can be formulated within a variational framework.
-Despite initial enthusiasm, standard VQAs face at least three critical challenges.
-First, optimization challenges, such as barren plateaus [9, 10, 11, 12, 13], where gradients vanish exponentially with circuit depth or system size, severely limit trainability.
-Second, measurement costs scale poorly with both qubit count and circuit parameters, placing heavy demands on quantum resources.
-Third, even hardware-efficient circuit designs may require long depths to represent certain wavefunctions, but decoherence makes such computations unreliable in practice [14].
-Meanwhile, deep learning has demonstrated remarkable capabilities in learning complex functions and high-dimensional structures from large datasets.
-In fact, neural networks (NNs) have set new standards for computational efficiency in tasks like quantum simulation [15, 16, 17, 18, 19], which are also commonly associated with quantum computing.
-A natural question then arises: can quantum resources enhance NN-based variational methods and help push the computational boundary of what is achievable with current and near-future hardware?
-Previous works have extensively investigated and documented that NNs can efficiently model the wavefunction amplitudes (i.e., essentially the probability distribution), but struggle to capture intricate sign or phase structures in many-body systems [20, 21, 22, 23, 19, 3].
-On the other hand, modeling the phase structure of an arbitrary wavefunction can be well approximated with some variations of parameterized Instantaneous Quantum Polynomial (IQP) circuit [24, 25, 26, 27, 28, 29].
-This observation motivates one way to combine PQC with NN to jointly model complex wavefunction within a hybrid variational framework.
-Importantly, such models may exhibit potential quantum advantage if continuous variable IQP (CV-IQP) operations exist for a single iteration [30, 31].
-Numerous efforts have explored the combination of quantum computing with deep learning in various settings [32, 33, 34].
-An especially relevant example is the variational quantum-neural hybrid eigensolver (VQNHE) [35] algorithm, previously developed by one of us and collaborators.
-This method uses a neural network to post-process quantum states prepared by a shallow circuit, forming a hybrid quantum-neural ansatz that efficiently expands the accessible quantum state space.
-However, VQNHE still faces limitations in trainability and incurs non-trivial measurement overhead.
-In light of these developments and challenges, we revisit how best to harness both quantum and classical resources within a variational framework, one that maximizes hybrid advantages while mitigating their downsides.
-To this end, we propose sign-VQNHE (sVQNHE), a hybrid algorithm that comprehensively assuages all the major challenges facing mainstream VQAs: expressiveness, measurement cost, trainability, and robustness (see Table 1).
-In short, the key innovations of sVQNHE lie in its principled separation of quantum and classical roles: the diagonal layers of the quantum circuit are tasked to model the phase structure (with much reduced measurement costs), while a non-negative neural network only has to learn the amplitude (or effectively, the probability) distribution for improved numerical stability.
-These components are coupled through an iterative amplitude-transfer mechanism that maps the neural state onto a shallow and classically simulatable quantum layer.
-Compared to the original VQNHE, this re-design of the hybrid quantum-neural framework not only optimizes expressivity with shallow circuits and substantially improves the measurement costs and the gradient-based convergence rate, but also inherently mitigates sampling noise and decoherence, yielding more robust solutions under realistic conditions (see Results).
-During the iterative, bidirectional feedback optimization loop (to train the PQC and NN parameters), if a single diagonal quantum layer cannot fully capture the target wavefunction’s sign structure, we extend the circuit depth by adding a layer of classically simulatable gates along with a new diagonal layer.
-The classical gate layer approximates the amplitude function learned by the NN by solving a tractable optimization problem, allowing the NN to guide amplitude refinement and assist in layer-wise circuit optimization without encountering barren plateaus.
-Importantly, because only the parameters of commuting diagonal gates are optimized on the quantum device, the required number of quantum measurements is dramatically reduced compared to the conventional non-commuting circuit ansatzes.
-In practice, at the last layer of the iterative, bidirectional feedback optimization procedure, one can choose to relax the constraint and allow the neural network to output complex values, thereby enhancing the expressiveness of the hybrid quantum-neural ansatz.
-Finally, our new algorithm improves the convergence behavior of the original VQNHE by incorporating a stable layer-by-layer training strategy that suppresses numerical instabilities.
-This newly proposed iterative and bidirectional feedback loop enables efficient training, expressive state representations, and significant reductions in circuit depth and measurement cost.
-Our numerical experiments validate these advantages.
-For instance, in the 6-qubit J1–J2 model, sVQNHE reduces mean absolute error by $98.9\%$ and variance by $99.6\%$ relative to a baseline NN, while achieving nearly 19× faster convergence compared to a hardware-efficient VQE using similar quantum resources.
-In solving the MaxCut problem on 45-node Erdős–Rényi graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$ over a brickwork ansatz VQE baseline [36].
-In summary, sVQNHE offers a scalable and practical path forward for tackling complex many-body quantum simulations and large-scale combinatorial optimization problems by fully leveraging the synergy between classical and quantum computation.
-As both quantum hardware and neural network models continue to advance, we anticipate that hybrid frameworks like sVQNHE will play an increasingly important role in achieving real-world quantum utility.
-Challenge Problem Description sVQNHE Solution
-Representation NN-based Variational Monte Carlo struggles to model complex phase structure Delegate the quantum circuit to learn phase structure
-Measurement
-VQE: $\mathcal{O}(2dn^{k^{\prime}})$, commuting gates compress groups to $\mathcal{O}(n^{k^{\prime}})$ ($k^{\prime}<k$)
-VQNHE: $\mathcal{O}(2dn^{k})$, $n$ qubit, $\mathcal{O}(n^{k})$ operators, $d$ parameter
-Commuting gates enable simultaneous measurements,
-reducing cost to $\mathcal{O}(n^{k})$
-Optimization
-VQE: barren plateau in deep circuits
-VQNHE: potential numerical instability
-Neural-guided layer-wise optimization with gradual
-amplitude transfer
-Noise
-VQE: affected by hardware noise and finite sampling noise
-VQNHE: NN mitigates errors but amplifies sampling noise
-Gradual amplitude transfer reduces sampling noise,
-NN mitigates errors
-Table 1: Key challenges in existing algorithms and sVQNHE’s mitigation strategies
+Hybrid quantum–classical approaches represent a promising strategy for extracting computational value from noisy intermediate-scale quantum (NISQ) and the subsequent early fault-tolerant quantum computing (Early FTQC) devices[preskill2018quantum, katabarwa2024early, nanda2024explorative].
+Central to this paradigm is the idea of variational wavefunction learning, in which a parametrized ansatz is iteratively optimized to minimize a task-specific objective, such as ground-state energy or a combinatorial cost function[peruzzo2014variational, mcardle2020quantum, cerezo2021variational].
+Despite substantial progress across physics, optimization, and quantum machine learning, current variational quantum algorithms (VQAs) scale poorly in the presence of noise and sampling constraints[cerezo2021variational, fedorov2022vqe].
+In practice, training often becomes unstable as system size grows, the gradient signal rapidly diminishes, and measurement cost increases with circuit complexity[cerezo2021cost, grant2019initialization, holmes2021barren, holmes2022connecting, wang2021noise, ilin2024dissipative].
+These failures stem from a profoundly imbalanced division of labor between quantum and classical resources.
+In most contemporary VQA architectures, from hardware-efficient layers to QAOA-style circuits, the quantum processor is often tasked to optimize a generic ansatz state with both the probability amplitudes and the delicate phase/sign structure to be tuned simultaneously.
+Capturing realistic amplitude distributions typically demands highly entangling gates and considerable circuit depth[grimsley2019adaptive, chen2020variational], whereas correctly learning the phase is extraordinarily fragile under hardware noise[cerezo2021cost, wang2021noise, ragone2024lie].
+As a result, improvements in one aspect often degrade the other, driving the well-documented trade-offs among circuit depth, trainability, sample complexity, and noise sensitivity.
+A largely unexplored yet critical question is therefore how to intelligently partition amplitude and phase learning between classical and quantum resources.
+Suppose the amplitude landscape can be captured classically[bluvstein2021controlling, lange2024architectures, pfau2024accurate, hermann2020deep, carleo2017solving] while reserving quantum resources for the inherently hard phase structure[westerhout2023many, park2025complexity, pagano2020quantum, leontica2024exploring, diez2024connection, szabo2020neural, westerhout2020generalization, nomura2017restricted, cai2018approximating, carleo2017solving, preskill2018quantum].
+In that case, the quantum circuit may remain shallow, measurement-efficient, and less affected by noise, without sacrificing overall expressivity.
+However, realizing such a division of labor poses its own challenges.
+The wavefunction representation must remain valid, optimization must be stable across heterogeneous parameter spaces, and the classical component must not eliminate the need for quantum computation altogether.
+Unlike classical variational quantum Monte Carlo methods, where amplitude-phase decomposition is straightforward, integrating quantum circuits introduces unique technical hurdles, such as ensuring non-commutative layers do not entangle the separated components and maintaining scalability under hardware constraints.
+In this work, we show that such a separation is not only possible but offers clear scaling advantages.
+We introduce sVQNHE, a variational framework in which a neural network (NN) models the amplitude distribution and a shallow diagonal quantum circuit learns the phase (sign) structure through neural-guided layer-wise optimization.
+To enable stacking multiple phase layers without collapse, since diagonal two-qubit gates alone support only quadratic interactions, we insert shallow, classically simulatable non-diagonal gates (e.g., a layer of $R_{y}$ gates) via a neural-guided amplitude transfer: the NN’s learned distribution is iteratively approximated onto these gates through classical simulations, preserving the decomposition by isolating amplitude uploads from phase encoding.
+This targeted quantum phase learning ensures that (i) dramatically reduced measurement cost due to commuting Pauli terms in the circuit, (ii) a well-conditioned optimization landscape that mitigates barren plateaus, (iii) significantly improved noise resilience from controlled circuit depth and NN mitigation, and (iv) more robust shot-based estimation of observables compared to the other non-unitary hybrid quantum-neural method.
+Notably, this design does not diminish quantum expressivity—neural amplitude modeling and quantum phase learning are jointly trained to reproduce nontrivial many-body states and combinatorial structure, representing a key innovation that bridges classical deep learning’s function approximation power with quantum hardware’s interference and entangling capabilities.
+We demonstrate sVQNHE’s effectiveness across strongly correlated quantum many-body systems and large-scale combinatorial optimization.
+On the 6-qubit frustrated J1-J2 model, sVQNHE reduces mean absolute energy error by $98.9\%$ and variance by $99.6\%$ versus the classical NN baselines, converging $\sim 19\times$ faster than hardware-efficient VQE.
+It also resolves molecular ground states such as the $H_{2}O$ molecule and demonstrates robustness under depolarizing noise and finite sampling.
+In the Maximum Clique problem (135 vertices, p=0.5), sVQNHE using only 10 qubits via Pauli-correlation encoding, substantially outperforms greedy heuristics, and a recently proposed Free-Energy Machine [shen2025free] (FEM), thereby confirming the non-trivial contribution of its quantum phase-learning module in highly constrained, rugged landscapes.
+Furthermore, on a MaxCut instance with 1,485 vertices, it achieves the same performance as the state-of-the-art, classical FEM using only 12 qubits.
+These results highlight sVQNHE’s enhanced generalization across many-body physics and diverse combinatorial optimization challenges, including problems where classical solvers like FEM face limitations due to mean-field approximations.
+This provides evidence that explicit quantum phase learning, tightly coupled with classical amplitude modeling, constitutes a principled paradigm for scalable hybrid algorithms.
+By optimizing the co-learning strategy on the amplitude and phase at the representation level, sVQNHE offers a pathway toward reliable near-term quantum utility in quantum chemistry, materials science, and optimization, enabling NISQ-era devices to deliver meaningful performance before the arrival of full fault tolerance.
 Results
 Theoretical Framework of sVQNHE
-Figure 1: (1) The algorithm seamlessly integrates quantum circuits with neural networks.
+Figure 1: (a) The algorithm seamlessly integrates quantum circuits with neural networks.
 The quantum circuits interleaves diagonal layers ($W_{1}\cdots W_{n}$) with shallow, simultable layers ($G_{1}\cdots G_{n}$).
@@
 Leveraging the commuting properties of the diagonal layers, the number of measurements needed to compute gradients for the $d$ parameters of $W$ drops from $\mathcal{O}(2d)$ to $\mathcal{O}(1)$.
-(2) Illustration of sVQNHE’s advantages: the star represents the optimal solution, the green circle represents the initial solution space that is easy to explore by the NN, the orange region represents the solution space accessible to the PQC, and the black circle represents the coverage of the hybrid quantum-classical scheme.
-Among them, (2a) describes the suboptimal situation, that is, PQC cannot connect the classical region with the optimal solution, and the hybrid scheme is difficult to cover the optimal solution; (2b) shows the potential quantum advantage; (2c) describes the transition strategy from (2a) to (2b) through parameter adjustment.
-(3) Illustration of the energy optimization process using the hierarchical optimization scheme.
-When entering the next layer, the local optimal solution will be jumped out due to the newly introduced parameters.
+(b) Illustration of sVQNHE’s advantages: The green dashed circle represents the coverage of the hybrid quantum-classical scheme.
+The star represents the optimal solution.
+(b1) describes the suboptimal situation, that is, the PQC cannot connect the solution region with the optimal solution; (b2) describes the potential quantum advantage.
+Similar to (b1), but the star is now inside the solution space, indicating successful coverage.
+(b3) Adjustable ansatz shows a transition from (b1) to (b2) via parameter adjustment, including gradient shading to indicate the evolution.
+(c) Illustration of the energy optimization process using the hierarchical optimization scheme.
+When entering the next layer, the local optimal solution will jump out due to the newly introduced parameters.
 We introduce sVQNHE, a hybrid quantum-classical variational algorithm designed to address key limitations of conventional VQAs, including high measurement cost, barren plateaus, limited expressive power with shallow circuits, and noise-induced errors.
-This framework builds on the original VQNHE approach by introducing a layer-wise, phase-amplitude decoupling strategy[24] (see Fig.1) that enhances trainability, scalability, and sampling robustness.
+This framework builds on the original VQNHE approach by introducing a layer-wise, phase-amplitude decoupling strategy[westerhout2023many] (see Figure 1) that enhances trainability, scalability, and sampling robustness.
 To elucidate the algorithm, we begin by summarizing the redesign of the quantum circuit and classical NN ansatzes. The overall (unnormalized) ansatz takes the form:
-|$\ket{\psi}=F\prod_{i=0}^{L-1}W_{i}G_{i}\ket{0},$ (1)
-where $F$ denotes a post-processing operator defined by a classical NN.
+[…]
+where $F_{L}$ denotes a post-processing operator defined by a classical NN.
 The following paragraphs explain the design of the hybrid ansatz and the associated bidirectional optimization strategy.
-Finally, the algorithm is summarized in Alg.
+Finally, the algorithm is recapped in Alg.
 1.
-Quantum Circuit for Phase Structure Learning.
-The quantum circuit ansatz consists of alternating layers of parameterized gates:$\left(G_{i},W_{i}\right)$.
-$G_{i}$ are classically simulatable and non-diagonal layers (e.g., $R_{y}$ gates), primarily responsible for shaping the amplitude structure.
-$G_{1}$ is usually taken to be a layer of Hadamard gates to create a uniform superposition.
-$W_{i}$ are diagonal gates (e.g., $R_{z}$ and $R_{zz}$) designed to encode the phase structure of the wavefunction.
-These layers are added iteratively when greater expressivity is needed, and their parameters are optimized in a layer-wise fashion.
+[…]
 Classical Neural Network for Amplitude Modeling.
-The classical neural network is constrained to produce non-negative outputs, and is used to model the amplitude $F=\sum_{s}f(s)\ket{s}\bra{s}$ | of the quantum state in the computational basis.
+The classical neural network is constrained to produce non-negative outputs, and is used to model the amplitude $|F=\sum_{s}f(s)\ket{s}\bra{s}$ of the quantum state in the computational basis.
 With the phase information delegated to the quantum circuit, the NN is focused on learning the probability distribution.
-Prior works have shown that classical neural networks are highly effective at this task, particularly for systems with smooth or structured amplitude landscapes [21, 24, 20].
+Previous works have shown that classical neural networks are highly effective in this task, particularly for systems with smooth or structured amplitude landscapes [westerhout2020generalization, westerhout2023many, szabo2020neural], relevant for many physical applications.
 Bidirectional Feedback Loop: Decoupling Amplitude and Phase.
-At each iteration $i$, a new block of paramterized quantum layers $(G_{i},W_{i})$ is added to the circuit.
+At each iteration $i$, a new block of parametrized quantum layers $(G_{i},W_{i})$ is added to the circuit.
 The process follows two key steps.
@@
 This ensures that $G_{i}$ inherits the amplitude structure previously learned by the NN.
-(2) Hybrid Optimization: Once $G_{i}$ is initialized, the neural operator $F$ is reset and jointly optimized with the parameters of the new diagonal gate layer $W_{i}$, as done in the original VQNHE algorithm.
+(2) Hybrid Optimization: Once $G_{i}$ has learn an approximation of $F$ and is loaded onto the quantum circuit, the neural operator $F$ is reset and jointly optimized with the parameters of the new diagonal gate layer $W_{i}$ by minimizing a loss function, such as the energy of a given Hamiltonian.
 This hybrid optimization simultaneously refines both the amplitude ($F$) and sign structure ($W_{i}$) in a tightly coupled quantum–classical feedback loop.
@@
 3:Initialization
-4:Initialize quantum state |$\ket{\psi_{0}}\leftarrow\ket{0}^{\otimes n}$
+[…]
 5:Initialize classical neural network parameters
-6:Define diagonal operator $F_{1}\leftarrow\sum_{s}f(s)\ket{s}\bra{s}$ | $\triangleright$ $f(s)$: NN-predicted amplitude
-7:Set $G_{1}\leftarrow H^{\otimes n}$triangleright Uniform superposition
+6:Define diagonal operator $|F_{1}\leftarrow\sum_{s}f(s)\ket{s}\bra{s}$ $\triangleright$ $f(s)$: NN-predicted amplitude
+7:Set $G_{1}\leftarrow H^{\otimes n}$ $\triangleright$ Uniform superposition
 8:Joint Optimization
-[…]
+9:while $\neg\text{converged}$ do $\triangleright$ Global convergence
 10:  for $l\leftarrow 1$ to $L$ do $\triangleright$ Layer-wise optimization
 11:   if $l>1$ then
-12:          Update $\theta^{g_{l}}$ to minimize $\|G_{l}-F_{l-1}\|$triangleright Gradual Transfer Mechanism, see Method for detail
+12:    Update $\theta^{g_{l}}$ to minimize $\|G_{l}-F_{l-1}\|$ $\triangleright$ Gradual Transfer Mechanism, see Method for detail
 13:   end if
-14:       $E_{l}\leftarrow 0$triangleright Reset layer energy
+14:   $E_{l}\leftarrow 0$ $\triangleright$ Reset layer energy
 15:Hybrid Energy Estimation
 16:   repeat
-17:          Apply operator: |$\ket{\psi_{l}}\leftarrow W_{l}(\theta^{w_{l}})G_{l}(\theta^{g_{l}})\ket{\psi_{l-1}}$
-18:          Construct hybrid state: |$\ket{\psi_{l,f}}\leftarrow F_{l}\ket{\psi_{l}}$
+[…]
+[…]
 19:    for each Pauli term $P_{i}\in H$ do
-20:             Compute $E_{P_{i}}\leftarrow\dfrac{\bra{\psi_{l,f}}P_{i}\ket{\psi_{l,f}}}{\innerproduct{\psi_{l,f}}{\psi_{l,f}}}$
+[…]
 21:      Accumulate energy: $E_{l}\leftarrow E_{l}+c_{i}\cdot E_{P_{i}}$
@@
 23:    Update NN parameters for $F_{l}$
-24:          Update parameters: $\theta^{w_{l}}$triangleright Measurement Strategy, see Method for detail
-25:       until $|\Delta E_{l}|<\epsilon_{\text{conv}}$triangleright Energy convergence
+24:    Update parameters: $\theta^{w_{l}}$ $\triangleright$ Measurement Strategy, see Method for detail
+25:   until $|\Delta E_{l}|<\epsilon_{\text{conv}}$ $\triangleright$ Energy convergence
 26:  end for
@@
 28:end while
-29:return $E_{L}$triangleright Final optimized energy
-Adaptive Hybrid Ansatz with Controlled Expressivity
-Conventional variational algorithms, such as the VQE and the QAOA, typically rely on predefined circuit ansatzes [37, 38].
-These ansatzes, often inspired by physical principles like Trotterized dynamics in QAOA, impose a static architecture that can limit their effectiveness.
-For example, if the ansatz fails to capture the complex entanglement or sign structure of the target ground state, the optimization may converge to suboptimal local minima.
-To address this, adaptive methods like adapt-VQE [39, 10] iteratively construct the ansatz by selecting gates from a predefined pool.
-However, these approaches encounter significant hurdles: the gate pool may not include operations capable of generating sufficient entanglement, thus constraining expressiveness, and the iterative gate selection process requires a large number of quantum measurements, making it impractical for large-scale systems.
-In contrast, VQNHE and sVQNHE employ a hybrid quantum-classical ansatz that effectively leverages both quantum and classical resources to represent a quantum state.
-This approach significantly expands the accessible quantum states with shallow quantum circuits.
-Nevertheless, the original VQNHE still assumes a fixed ansatz approach, which could clearly be optimized to further reduce circuit depth.
-To this end, sVQNHE adopts a neural-guided framework to systematically extend the circuit’s depth along with controlled expressiveness.
-Its shallow quantum layers $G_{l}$ are sequentially determined under the gradual guidance of a classical neural network.
-A key feature of sVQNHE is its ability to achieve high expressiveness and stable convergence even with simple layers of $G_{i}$.
-For instance, take $G_{i}$ to be a layer of $R_{y}$ rotations on each qubit, can already enhannce the ansatz’s flexibility.
-This capability is rigorously supported by Theorem 1 (Expressiveness of the Quantum Circuit Ansatz), which analyzes the Lie algebra generated by specific sets of operators [40, 41, 12, 42, 43].
-Consider the generator set $\mathcal{G}_{1}=\{Z_{D_{2j}}\mid j=1,2,\ldots,\lfloor\frac{m}{2}\rfloor\}\cup\{Y_{i}\mid i=1,2,\ldots,n\}$, where $Y_{i}$ are Pauli Y operators on individual qubits, and $Z_{D_{2j}}$ are Pauli Z operators on specific qubit subsets.
-The theorem proves that the corresponding Lie algebra is $\mathfrak{g_{1}}=\mathfrak{su}(2^{n-1})\oplus\mathfrak{su}(2^{n-1})$, a highly expressive structure with a dimension far exceeding that of alternative generator sets (e.g., $\mathcal{G}_{2}$).
-This large Lie algebra indicates that even layers primarily composed of $R_{y}$ gates, can explore a vast state space, enhancing the ansatz’s ability to approximate complex quantum states efficiently.
-Moreover, the design of the quantum layers $G_{l}$ in sVQNHE is guided by structural constraints that ensure computational efficiency and scalability.
-Proposition 1 (Commuting Block Structure) provides a critical insight: if a block diagonal matrix $M=M_{1,1}\oplus\cdots\oplus M_{m,m}$ has distinct eigenvalues across its blocks, any matrix commuting with $M$ must preserve the same block structure.
-In the context of sVQNHE, this proposition informs the construction of $G_{l}$ layers.
-By respecting this commuting block structure, sVQNHE can increase circuit depth in a controlled manner, isolating commuting blocks (e.g., $W_{i}$).
-Unlike rigid ansatzes (e.g., QAOA’s Trotterized structure) or gate-pool-dependent methods (e.g., adapt-VQE), sVQNHE avoids both architectural rigidity and measurement bottlenecks, achieves higher expressivity with less quantum resource demands.
-Efficient Measurement and Trainability via Diagonal Gates and Neural Guidance
+29:return $E_{L}$ $\triangleright$ Final optimized energy
+Efficient Allocation of Quantum Resources
+Traditional variational algorithms, such as the Variational Quantum Eigensolver (VQE) and the Quantum Approximate Optimization Algorithm (QAOA), typically employ fixed circuit ansatzes derived from physical intuitions like Trotterized time evolution[aaronson2009need, mcclean2021low].
+These static architectures can limit performance when the ansatz fails to capture the entanglement or correlations required for the target state, often trapping optimization in local minima.
+Adaptive strategies like adapt-VQE attempt to address this by iteratively selecting gates from a predefined pool, but they remain constrained by the expressiveness of the pool and incur high measurement overhead due to repeated gate evaluations[tang2021qubit, grant2019initialization].
+In contrast, sVQNHE employs a neural-guided adaptive hybrid ansatz (as defined in Equation (1)) that achieves an alternative and (potentially) more efficient allocation of quantum resources.
+The ansatz alternates between shallow and non-diagonal $G_{i}$ (e.g., $R_{y}$ rotations) and diagonal multi-qubit phase-encoding layers $W_{i}$ (e.g., $R_{z}$ and $R_{zz}$).
+Crucially, circuit growth proceeds in a layer-wise manner under the guidance of the classical neural post-processing module $F$.
+At each stage of optimization, the neural network receives the current output distribution from the quantum circuit and provides feedback that informs whether additional layers are required and, if so, how they should be structured to most effectively refine the approximation.
+This amplitude-guided, progressive deepening ensures that quantum resources are deployed selectively and only where necessary to capture residual discrepancies in the target distribution, avoiding unnecessary depth or parameter overhead.
+Complementing this adaptive construction is a measurement strategy that exploits the commuting structure within each layer: the diagonal $W_{i}$ layers are fully commuting with extremely low measurement costs, while the parameters in shallow, simulatable $G_{i}$ layers are optimized classically without any measurement costs.
+As a result, gradient estimation requires substantially fewer quantum shots compared with the standard VQE algorithm employing the hardware-efficient or other general, entangling ansatzes, thereby reducing the overall measurement burden and enhancing scalability on devices with limited shot capacity.
+A detailed discussion of the measurement-efficient gradient protocol is provided in the Measurement Reduction part.
+Even with relatively simple non-diagonal layers based on $R_{y}$ rotations, the chosen generator set $\{ZD_{2j},Y_{i}\}$ spans a Lie algebra of dimension $\mathfrak{su}(2^{n-1})\oplus\mathfrak{su}(2^{n-1})$ (Theorem 1), enabling the approximation of highly entangled states using only shallow circuit depth.
+Furthermore, the incorporation of neural post-processing helps maintain polynomial gradient variance in the loss landscape, $\mathrm{Var}[\partial_{\theta_{k}}\mathcal{L}]\in\Omega(1/\mathrm{poly}(n))$, thereby suppressing the emergence of barren plateaus during adaptive growth (Theorem 2).
+Proposition 1 additionally informs the design of non-commuting layer sequences, ensuring that added depth contributes meaningfully to expressiveness.
+Consequently, sVQNHE dynamically assembles shallow yet highly expressive circuits whose reachable Hilbert space expands in a targeted, data-driven fashion until the ground-state wavefunction or desired combinatorial solution distribution is adequately represented (illustrated schematically in Figure 1(b3)).
+Empirical results across frustrated spin models (such as J1–J2 chains) and large-scale combinatorial optimization instances (including MaxCut and Maximum Clique under qubit-efficient encodings) demonstrate that sVQNHE consistently achieves high-fidelity output distributions using fewer quantum layers and markedly lower measurement costs than fixed-depth variational baselines, non-adaptive hybrid methods, or purely classical neural approaches, even for distributions that are classically difficult to sample due to strong entanglement and complex sign structure [troyer2005computational].
+In summary, by integrating neural-directed adaptive growth, measurement-frugal layer designs, and Lie-algebra-enhanced shallow-circuit expressiveness, sVQNHE reallocates quantum resources far more efficiently than traditional fixed or purely adaptive variational frameworks.
+This resource-aware construction positions the method as a robust candidate for realizing practical quantum utility on near-term noisy quantum devices, where both qubit count and measurement budget remain severely constrained.
+Reducing Measurement Costs and Enhancing Trainability
 Traditional variational quantum algorithms face two intertwined challenges that limit scalability: high measurement costs and the emergence of barren plateaus.
@@
 The sVQNHE framework addresses these limitations through two synergistic innovations.
-First, it employs commuting diagonal gates, such as $R_{z}$ and $R_{zz}$ rotations, which enable simultaneous measurement of expectation values in one basis.
+First, it employs commuting diagonal gates, such as $R_{z}$ and $R_{zz}$ rotations, which enable simultaneous measurement of all relevant expectation values in one basis.
 This compresses the per-operator measurement cost to $\mathcal{O}(1)$, reducing the total overhead to $\mathcal{O}(\text{poly}(n))$ for $\text{poly}(n)$ operators (see details in the method section).
 This stands in stark contrast to the multiplicative $\mathcal{O}(2d\cdot\text{poly}(n))$ scaling of traditional VQNHE and conventional VQAs, allowing sVQNHE to tackle large-scale problems efficiently.
-Secondly, our framework incorporates layer-wise neural guidance to decompose the optimization process into manageable subproblems, addressing a critical limitation of traditional layer-wise VQAs[44].
+Secondly, our framework incorporates layer-wise neural guidance to decompose the optimization process into manageable subproblems, addressing a critical limitation of traditional layer-wise VQAs[liu2022layer].
 In conventional layer-wise VQAs, the absence of a global perspective on the optimization landscape renders sequential layer optimization prone to converging at local minima.
@@
 This neural guidance not only helps avoid suboptimal local minima but also focuses parameter updates on one specific block each time, such as $W_{l}$ at the $l$-th iteration.
-This approach ensures that gradients remain measurable even in deep circuits, since the effective parameter space at each optimization step can always be restricted to a minuscule fraction of the full deep circuit.
+This approach, supported by the benign landscape of Theorem 2, restricts the effective parameter space per iteration (e.g., block $W_{l}$), preventing convergence to local minima and maintaining scalability.
 Together, these strategies circumvent the dual bottlenecks of measurement overhead and barren plateaus, establishing a scalable pathway for variational quantum algorithms in practical applications.
@@
 The variance of the energy estimate $\hat{E}_{P_{i}}$ is governed by:
-$\text{Var}(\hat{E}_{P_{i}})=\text{Var}\left(\bra{\psi_{L,f}}P_{i}\ket{\psi_{L,f}}\right)\cdot\text{Var}(g^{-1}(s,L)),$ (2)
+[…]
+[…]
 where $g(s,L)=\frac{1}{n_{s}}\sum_{s_{i}}f^{2}(s_{i})$, and $s_{1},\ldots,s_{n_{s}}$ are i.i.d.
@@
 $\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left(f^{2}(s)\right)\xrightarrow{L\uparrow}0.$ (3)
-For sufficiently large $L$, $\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left(f^{2}(s)\right)<1$. By the uniform convergence of the series,
-$\text{Var}(g^{-1}(s,L))=\sum_{i}\mu_{i}\text{Var}^{i}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left[f^{2}(s)\right],$ (4)
-with $\mu_{i}>0$.
-This variance asymptotically approaches zero as $L\to\infty$.
-In our numerical experiment, as illustrated in Figures 3 and 4, the coefficient of variation of energy $\text{CV}\propto\sqrt{Var(g^{-1}(s,L))}$ decays with $L$, suggesting that the algorithm has better noise immunity under actual sampling constraints as $L$ increase.
-In summary, we eliminate a source of numerical instability (which subsequently affects the convergence rate) of the original VQNHE.
+For sufficiently large $L$, the variance of $f^{2}(s)$ becomes small ($\text{Var}(f^{2}(s))<1$), which implies that $g(s,L)$ concentrates tightly around its mean $\mathbb{E}[g(s,L)]=\mathbb{E}\left[f^{2}(s)\right]$.
+Under this concentration condition, the variance of the normalization factor $g^{-1}(s,L)$ can be accurately approximated using the first-order delta method (also known as the linearization method):
+$\text{Var}\bigl(g^{-1}(s,L)\bigr)\approx\frac{1}{\mathbb{E}^{4}[g(s,L)]}\text{Var}(g(s,L))=\frac{1}{\mathbb{E}^{4}[g(s,L)]}\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\bigl[f^{2}(s)\bigr].$ (4)
+[…]
+In our numerical experiments, as illustrated in Figures 4(a) and (b), the coefficient of variation of the energy estimate $\text{CV}(\hat{E})\propto\sqrt{\text{Var}(g^{-1}(s,L))}$ decays with increasing $L$.
+This confirms that sVQNHE exhibits improved noise immunity and robustness under realistic finite-sampling constraints as the quantum circuit depth grows.
+In summary, we mitigate a source of numerical instability of the original VQNHE.
 Experimental Results
-To rigorously validate the accuracy and efficiency of our proposed algorithm for the ground-state preparation of a given Hamiltonian, we conducted systematic evaluations across four canonical manybody problems: the 1D J1-J2 model, the 2D Heisenberg model, the Transverse Field Ising Model (TFIM), and the Classical Ising model.
-These models were chosen to represent diverse physical systems with varying degrees of sign problem severity and ground-state complexity.
-We employed three methodologies, sVQNHE, VQE, and NNs, to benchmark their performance.
-Additionally, we integrated the encoding scheme proposed in Ref.
-[36] to demonstrate the superiority of sVQNHE over conventional VQAs (such as QAOA and VQE) in solving the MaxCut problem, a prototypical combinatorial optimization task.
-Our experiments included a 6-qubit 1D J1-J2 model with $J1/J2=0.6$, a configuration known to exhibit a highly non-trivial ground-state phase structure [45].
-This model serves as a representative case for systems with intricate sign signatures.
-We also tested a 9-qubit 2D Heisenberg model with periodic boundary conditions, which is known to incur severe sign-related challenges for numerical methods[46].
-For comparison, we analyzed a 9-qubit 1D TFIM (solvable via Jordan-Wigner transform) and two 1D Ising models (6 and 12 qubits).
-Both TFIM and Ising models we chose are Sign-problem-free systems.
-Nevertheless, they still present non-trivial challenges in different ways.
-For example, they are relevant to the qubit-formulated combinatorial optimization problems.
-We evaluated our algorithm under four scenarios to assess its theoretical advantages and practical robustness:
-  • Ideal Simulation: Quantified intrinsic accuracy, variability, and scalability under ideal conditions.
-  • Real-World Simulation: Assessed robustness under realistic quantum conditions, combining depolarization-based gate errors (two-qubit: $0.005$; single-qubit: $0.0013$) with bitstring-sampling-based estimation of observables and gradients.
-  • Scalability Under Finite Sampling: Demonstrated that sVQNHE’s advantage over existing methods[36] can be sustained with system size.
-We emphasize that the results of all the so-called "non-ideal" experiments are based on the original sampling data (of bitstrings) from noisy quantum circuits, without implementing any quantum error mitigation procedures.
-As sVQNHE possesses an inherent and robust error mitigating capability [47] due to the fact that the hybrid quantum-neural ansatz can be directly tuned with the classical NN module during the noisy optimization process.
-1. Ideal Simulation
-We conducted a systematic evaluation of the performance of sVQNHE and other methods (VQNHE, VQE, NN, QAOA) focusing on accuracy, variability, and measurement efficiency.
-All the calculations were done in statevector-type exact numerical computations.
-For consistency, sVQNHE and VQNHE utilized the same non-negative real neural network component (MLP), while the NN method employed a real MLP with identical parameters and structure.
-The quantum components of sVQNHE and VQNHE were our single-layer sign ansatz and a hardware-efficient ansatz with a similar number of parameters, respectively.
-In Table 2, these are denoted as sign-NN and hea-NN, respectively.
-The VQE method used a hardware-efficient ansatz with varying layers, such as HEA2 for two layers.
-Similarly, the QAOA method was tested with different layers, for instance, QAOA10 for ten layers.
-All methods initialized quantum circuit parameters using independent uniform distributions $U(-2\pi,2\pi)$, optimized for the same number of steps, and averaged the estimated energy from 20 random initial points.
-The baseline algorithm in Table 2 is NN.
-Simulating Frustrated Many-Body Systems
-For the 1D J1-J2 model, which features a non-trivial ground state, the sVQNHE method (sign-NN) demonstrated significant improvements over the NN method, the VQE method (HEA2), and the VQNHE method (hea-NN).
-Specifically, Table 2 shows that after 2000 optimization steps, the sign-NN method reduced the mean absolute error (MAE) by $98.9\%$ and the variance by $99.6\%$ for the 6-qubit J1-J2 model compared to the NN method.
-Additionally, Figure 2 illustrates the scaling results (from 6 qubits to 12 qubits) of sign-NN and NN with the same number of optimization steps to achieve the minimum error.
-The advantage of sign-NN increases with the number of qubits.
-In contrast, while hea-NN also reduced the MAE by $75.1\%$ compared to NN, its variance was larger than that of NN (variance: hea-NN > NN > sign-NN).
-Furthermore, due to the non-commutative structure of hea, the measurement amount for sign-NN was significantly smaller than that for hea-NN under similar parameter amounts.
-These three points—improved accuracy, reduced variance, and reduced measurement amount—demonstrate the superiority of sVQNHE over the original VQNHE algorithm.
-Regarding convergence speed, sign-NN significantly outperformed HEA2.
-To achieve the same error threshold, HEA2 required 12.6 times more optimization steps on average than sign-NN.
-In the most efficient cases, HEA2 needed 1676 steps to reach the same accuracy, while sign-NN only required 89 steps.
-Additionally, the probability of the sign-NN successfully reaching the target accuracy was 12 times higher than that of HEA2, highlighting its excellent reliability and efficiency (for details, please refer to the supplementary material).
-Similar results were observed in the Heisenberg model, further verifying the capability of sVQNHE to handle systems affected by the sign problem.
-Model Method $R_{MAE}$ $R_{Var}$
-6-qubit 1D J1-J2 sign-NN -0.989 -0.996
-hea-NN -0.751 1.955
-HEA2 -0.916 -0.132
-9-qubit 2D Heisenberg HEA1 -0.331 0.554
-hea-NN -0.119 33.793
-sign-NN -0.872 -0.832
-9-qubit 1D TFIM HEA1 -0.290 0.019
-sign-NN -0.640 -0.728
-9-qubit 1D Ising HEA1 0.129 0.909
-QAOA5 20.674 15.583
-QAOA10 1.205 11.835
-sign-NN -0.416 -0.455
-12-qubit 1D Ising HEA1 6.822 10.002
-QAOA10 77.012 80.404
-sign-NN -0.811 -0.979
+To systematically evaluate the proposed sVQNHE framework, we design a 4-stage benchmark suite that progressively demonstrates its key advantages: (1) its ability to overcome optimization challenges arising from complex phase structures in strongly frustrated quantum many-body systems, (2) its general performance improvement even in systems with trivial phase structures due to enhanced optimization landscapes and measurement efficiency, (3) its robustness and noise mitigation capabilities under real NISQ conditions (without any error mitigation), and (4) its potential to deliver near-term quantum utility in large-scale combinatorial optimization using a qubit-efficient encoding.
+All quantum simulations are performed without any form of quantum error mitigation.
+Noisy results are obtained directly from shot-based sampling of noisy quantum circuits.
+Unless otherwise stated, energy values are reported as relative errors with respect to the exact ground-state energy, and all results are averaged over multiple random initializations.
+1. Frustrated Quantum Many-Body Systems
+The experimental application of the sVQNHE framework to frustrated quantum spin systems serves as a critical validation of its advantages.
+These systems, characterized by highly complex ground-state phase structures and strong correlations, pose significant challenges to conventional variational quantum algorithms.
+Because it requires deep circuits to optimize both amplitude and phase components simultaneously, they often encounter optimization challenges, such as getting stuck in barren plateaus or local minima [mcclean2018barren, cao2025exploiting].
+By explicitly separating these tasks, delegating phase learning to a shallow quantum module while modeling the amplitude with classical neural networks, sVQNHE effectively mitigates the complex phase structures problem.
+The resulting faster and more stable convergence demonstrates that sVQNHE successfully alleviates key optimization dilemmas inherent to traditional VQEs, thereby achieving superior performance in strongly correlated systems compared to purely classical or purely quantum approaches and highlighting the framework’s potential as an efficient hybrid strategy for tackling complex quantum many-body problems.
+We focus on two paradigmatic models with highly non-trivial ground-state phase diagrams and strong sign oscillations in the computational basis to test:
+    the 1D J1–J2 chain with $J_{1}/J_{2}=0.6$ (in the Majumdar–Ghosh regime near the dimerized phase, known for its challenging phase structure [viteritti2022accuracy]),
+    the 2D square-lattice Heisenberg antiferromagnet with periodic boundary conditions (9 qubits, exhibiting severe phase complexity in numerical methods [stoudenmire2012studying]).
+We compare sVQNHE (sign-VQNHE) against four strong baselines:
+    pure classical neural network (NN) with real-valued MLP,
+    hardware-efficient VQE (HEA) with 1 or 2 layers,
+    original VQNHE using the same hardware-efficient quantum circuit and real-valued NN (denoted hea-NN),
+    QAOA with up to 10 layers.
+All methods are trained with the same number of optimization steps (2000) and evaluated over 20 random initializations.
+Performance is quantified using relative mean absolute error ($R_{\text{MAE}}$) and relative variance ($R_{\text{Var}}$) with respect to the pure NN baseline (see Methods for definitions; values closer to -1 indicate better performance).
+Results are summarized in Table 1.
+Metric 6q J1-J2 9q Heisenberg 9q TFIM 9q Ising 12q Ising
+sign-VQNHE hea-NN HEA2 HEA1 hea-NN sign-VQNHE HEA1 sign-VQNHE HEA1 QAOA5 QAOA10 sign-VQNHE HEA1 QAOA10 sign-VQNHE
+$R_{\text{MAE}}$ -0.989 -0.751 -0.916 -0.331 -0.119 -0.872 -0.290 -0.640 0.129 20.674 1.205 -0.416 6.822 77.012 -0.811
+$R_{\text{Var}}$ -0.996 1.955 -0.132 0.554 33.793 -0.832 0.019 -0.728 0.909 15.583 11.835 -0.455 10.002 80.404 -0.979
 Table 1: Performance Analysis of sVQNHE Algorithm.
 Here, the range of $R_{MAE}$ and $R_{Var}$ is $(-1,+\infty)$.
-The closer to $-1$, the better (see Equations (12) and (13) in the Method section for details).
+The closer to 1-1, the better (see Equations (LABEL:eq:R_MAE) and (LABEL:eq:R_var) in the Method section for details).
 The baseline algorithm NN model is a real Multi-Layer Perceptron (MLP).
@@
 The Sign Ansatz employs $R_{zz}$ gates to connect all edges of the corresponding physical model.
+For the 6-qubit J1-J2 chain, sVQNHE achieves dramatic improvements.
+Specifically, $R_{\text{MAE}}=-0.989$ implies $98.9\%$ error reduction vs.
+NN, and $R_{\text{Var}}=-0.996$ implies $99.6\%$ variance reduction vs.
+NN.
+In contrast, hea-NN only reduces MAE by $75.1\%$ while increasing variance compared to NN.
+Conventional 2-layer HEA (HEA2) reaches -0.916 in MAE but still lags far behind sVQNHE in both accuracy and stability.
+Moreover, sVQNHE converges significantly faster: it requires on average $12.6\times$ fewer steps than HEA2 to reach the same error threshold, and succeeds in reaching target accuracy $12\times$ more frequently.
+In the most efficient cases, HEA2 needed 1676 steps to reach the same accuracy, while sVQNHE only required 89 steps (for details, please refer to the supplementary material).
+Furthermore, due to the non-commutative structure of hardware-efficient ansatze, the measurement overhead for sVQNHE is significantly lower than for hea-NN under similar parameter counts.
+The scaling behavior is even more striking.
+Figure 2 shows the performance gap between sVQNHE and NN as system size increases from 6 to 12 qubits under a fixed optimization budget.
+The advantage of sVQNHE widens systematically with system size, demonstrating that the phase-focused architecture becomes increasingly valuable as correlation strength and phase complexity intensify.
 Figure 2: Scaling test for the 1D J1-J2 model.
-The x-axis shows the number of qubits, while the y-axis displays the average relative error from the sVQNHE method (sign-NN, blue line) and the NN method (NN, orange line) after 20 runs of 2000 steps each.
-The green line illustrates the difference in average relative error between the NN and sign-NN methods as the qubit count increases.
+The x-axis shows the number of qubits, while the y-axis displays the average relative error from the sVQNHE method (sign-VQNHE, blue line) and the NN method (NN, orange line) after 20 runs of 2000 steps each.
+The green line illustrates the difference in average relative error between the NN and sign-VQNHE methods as the qubit count increases.
 Shaded regions represent the relative error bar scaled by 0.5.
-Simulating Sign-Problem-Free Many-Body Systems
-Although the above results highlight the advantages of sVQNHE in sign-frustrated quantum models, the same architectural design is effective even in sign-free systems, where phase modeling is trivial for the ground state itself, but the ability to explore a broader Hilbert space (as defined by the PQC-generated phase strcutures) turns out to be beneficial for the optimization convergence in the numerical investigated presented in this subsection.
-More precisely, please see the last three rows of Table 2.
-In our numerical experiments on the Ising model, the sign-NN approach demonstrates substantial improvements over the NN in both error reduction and variance compression.
-At 9 qubits, sign-NN achieves a $41.6\%$ reduction in MAE compared to NN, indicating a significant enhancement in accuracy.
-This advantage becomes even more pronounced at 12 qubits, where sign-NN reduces MAE by $81.1\%$ relative to NN, showcasing its superior scalability with increasing system size.
-Similarly, in terms of variance, sign-NN outperforms NN by reducing variance by $45.5\%$ at 9 qubits.
-This gap widens dramatically at 12 qubits, with sign-NN achieving a $97.9\%$ reduction in variance compared to NN, highlighting its ability to deliver consistent and reliable results as the system scales.
-When benchmarked against the QAOA in the 9-qubit Ising model, sign-NN maintains its performance edge.
-A 5-layer QAOA configuration requires a comparable number of per-iteration measurements to sign-NN but yields a higher MAE, indicating lower accuracy.
-Extending QAOA to 10 layers improves its accuracy, yet it still falls short of sign-NN’s performance.
-This trend persists and becomes more evident at 12 qubits, where sign-NN’s scalability further amplifies its advantage over QAOA.
-The consistency of these results extends to the TFIM.
-The superior performance of sign-NN can be attributed to its carefully designed ansatz, optimized with neural network guidance.
-Unlike QAOA, which uses $R_{x}$ gates and constrains all $Z$- and $X$-related gates within a layer to share a single parameter, the Sign Ansatz employs $R_{zz}$ and $R_{y}$ gates with independent parameters.
-This design, as supported by Theorem 1, generates a higher-dimensional Lie algebra ($\mathfrak{g}_{1}$) compared to QAOA’s ($\mathfrak{g}_{2}$), enabling sign-NN to explore a broader unitary space through multi-qubit $Z$-phase adjustments.
-The phase-focusing architecture of the Sign Ansatz, which leverages multi-qubit $Z$-gates to refine phase information, allows for efficient navigation of high-dimensional solution spaces.
-When combined with neural network optimization, this approach achieves high accuracy and low variance at a reduced computational cost, making sVQNHE particularly well-suited for large-scale quantum optimization problems.
-2. Real-World Simulation
-Next, we consider how robustly sVQNHE performs when there exists both gate errors (for e.g.
-depolarization channels) and uncertainty induced by the finite sampling of bitstrings instead of statevector calculations as in the previous section.
-We evaluated sVQNHE against VQE, adopting two distinct optimization strategies, layered-VQE (layerwise optimization) and standard-VQE (joint optimization of all parameters), using an identical ansatz circuit.
-sVQNHE employed non-negative real neural networks, while both VQE methods utilized the same Sign Ansatz without NN augmentation.
-For all VQE schemes, the resource consumption was saved by measurement based on commuting groups.
-All simulations capped iterations at 400.
-Figure 3: Performance evaluation of sVQNHE and VQE variants under a 3-qubit 1D Heisenberg Hamiltonian.
-Energy (y-axis) versus iteration count (x-axis), with the ground state energy marked by a black dashed line.
-For sVQNHE and layered-VQE (layered optimization), blue and orange lines correspond to the first and second evolutionary layers, respectively, where each layer introduces new variational parameters.
-For standard-VQE (overall optimization), the lines reflect energy progression across sequential steps under fixed parameterization.
-For sVQNHE, coefficients of variation after stabilization (final 50 iterations per layer) are $cv_{1}=0.01618$ and $cv_{2}=0.015912$, indicating a reduction in variability in the second layer.
-Figure 4: Energy convergence of sVQNHE and VQE variants under a noisy 3-qubit 1D Heisenberg Hamiltonian.
+Similar qualitative behavior is observed on the 9-qubit 2D Heisenberg model ($R_{\text{MAE}}=-0.872$, $R_{\text{Var}}=-0.832$), confirming that sVQNHE consistently resolves the trainability barriers faced by both pure classical NN (which cannot model phases) and hardware-efficient variational circuits (which entangle amplitude and phase learning) in regimes with complex phase structures.
+2. Trivial Phase Structures Systems
+Having demonstrated sVQNHE’s strength in systems with complex phase structures, we next ask whether the method remains advantageous in systems where the ground-state phase is trivial (i.e., real-valued or easily captured without specialized phase modeling).
+We consider three representative cases without significant phase frustration:
+    1D Transverse Field Ising Model (TFIM, 9 qubits, exactly solvable via Jordan–Wigner transformation),
+    1D classical Ising chains (9 and 12 qubits, relevant to combinatorial optimization encodings),
+    electronic ground state of the water molecule ($H_{2}O$) in STO-3G basis (10 qubits after active-space reduction, introducing electronic correlation challenges).
+Table 1 shows that even without complex phases, sVQNHE systematically outperforms the pure NN baseline.
+Notably, the performance margin increases with system size, indicating improved scalability.
+When benchmarked against QAOA (5 or 10 layers), sVQNHE maintains its edge: QAOA requires more measurements per iteration and yields higher errors, even at deeper layers.
+The superior performance stems from sVQNHE’s phase-focusing architecture (using $R_{zz}$ and $R_{y}$ gates with independent parameters), which generates a higher-dimensional Lie algebra compared to QAOA (as per Theorem 1), enabling broader exploration of the unitary space.
+For the $H_{2}O$ molecule, we compare sVQNHE against standard 3-layer hardware-efficient VQE across multiple geometries (generated with PySCF, bond length $1.2$, bond angles $90^{\circ},120^{\circ}]$).
+Both methods use comparable total quantum parameter counts.
+As shown in Figure 3, sVQNHE consistently achieves: lower normalized energy error and higher wavefunction fidelity.
+Across the entire geometry scan, despite using half as many quantum parameters per iteration and benefiting from fully-commuting layers (leading to dramatically lower measurement cost).
+These results demonstrate that the division of the amplitude–phase task between quantum and classical resources with neural-guided optimization significantly improves trainability and solution quality even when the phase learning is not particularly challenging.
+The benefits arise from a smoother optimization landscape, reduced variance, and enhanced measurement efficiency, advantages that are broadly relevant across many problem classes.
+Figure 3: Comparison of normalized energy and fidelity for $H_{2}O$ molecules at various bond angles (angle).
+The x-axis labels denote different geometric configurations (r, angle), where (r) is the bond length.
+The top panel (a) presents the average normalized energy $E_{\mathrm{pred}}/E_{\mathrm{exact}}$ for each group, while the bottom panel (b) shows the average fidelity between the predicted and exact wavefunctions.
+Blue bars represent results from the VQE method, and orange bars correspond to the sVQNHE method.
+Error bars indicate 0.5 standard deviation over 5 independent runs.
+3. Robustness Under Noisy Conditions
+To assess the practical viability of sVQNHE in the NISQ regime, we evaluate its performance under two realistic sources of imperfection: (1) finite sampling shot noise from bitstring measurements, and (2) hardware gate and readout noise (modeled via depolarizing channels).
+All results are obtained directly from noisy circuit sampling without applying any quantum error mitigation techniques.
+We focus on the 3-qubit 1D Heisenberg Hamiltonian as a representative frustrated system.
+We compare sVQNHE against two VQE variants using the identical Sign Ansatz: layered optimization VQE (layerwise-VQE) and standard optimization VQE (standard-VQE).
+sVQNHE employs non-negative real-valued neural networks for amplitude correction, while both VQE baselines optimize only quantum parameters.
+All simulations are capped at 400 iterations, with resource-efficient commuting-group measurements for VQE.
+Figure 4(a) shows convergence curves (energy vs.
+iteration) under finite sampling (ideal noise-free circuits, limited shots).
+sVQNHE converges faster and more stably than both VQE variants.
+Notably, layered optimization fails for pure VQE because the 1-layer Sign Ansatz leads to premature trapping in poor local minima.
+In contrast, sVQNHE successfully leverages layered growth by delegating initial amplitude modeling to the NN, enabling effective progressive refinement.
+Post-convergence, the coefficient of variation (CV) in the second layer is lower than in the first ($CV_{1}=0.01618\to CV_{2}=0.015912$), confirming that the gradual transfer mechanism enhances stability under sampling variance.
+(a) Performance evaluation under finite sampling.
+(b) Performance evaluation under noisy conditions.
+Figure 4: Comparative performance of sVQNHE and VQE variants under a 3-qubit 1D Heisenberg Hamiltonian.
+(a) Finite sampling (1200 shots); (b) Noisy conditions (1200 shots, two-qubit: $0.005$; single-qubit: $0.0013$).
 Iteration count (x-axis) versus energy (y-axis).
@@
 For layered optimization (sVQNHE, layered-VQE), blue and orange lines denote first- and second-layer parameter updates, respectively; standard-VQE (overall optimization) reflects sequential parameter refinement.
-For sVQNHE, coefficients of variation after stabilization (final 50 iterations per layer) are $cv_{1}=0.01534$ and $cv_{2}=0.01487$, indicating a reduction in variability in the second layer.
-As shown in Figures 3 and 4, sVQNHE consistently achieved faster convergence than VQE, both with and without device noise.
-While layered optimization inherently reduces computational resource demands, mitigates noise sensitivity via shallower circuits, and often accelerates convergence, it is not applicable to the general VQE framework due to premature convergence to local optimal solutions.
-Specifically, the 1-layer Sign Ansatz fails as a viable initial parameterization for VQE’s layered optimization, as evidenced by the divergent performance of layer-VQE and standard VQE in our tests.
-However, integrating NNs into the Sign Ansatz (as in sVQNHE) circumvents this limitation, enabling effective layered optimization by delegating partial modeling duty to the neural network.
-Post-convergence analysis of sVQNHE revealed that the coefficient of variation (CV) for the second optimization layer remained lower than that of the first layer across all noise conditions.
-This validates the efficacy of our gradual transfer mechanism, which systematically refines solutions while maintaining stability under noise.
-3. Scaling Under Sampling Errors
-Finally, we investigate whether the observed advantage of sVQNHE can be sustained as the system size increases under the real-world setting of extracting PQC state by sampling bitstrings on quantum circuits.
-An example is provided in Figure 5.
-We evaluate the solution quality and measurement cost of the MaxCut problem on Erdős–Rényi random graphs with an edge connection probability of $0.3$ as the graph size increases.
-To establish this advantage, we compare sVQNHE against the Sign Ansatz VQE (sign-VQE) and the recently proposed brickwork-VQE [36], all evaluated under the same polynomial encoding model.
-This model encodes the $m$ binary variables of the MaxCut problem into the Pauli correlations across $k$ qubits, enabling the representation of $m=O(n^{k})$ variables with only $n$ qubits, where $k>1$ is an integer.
-Unlike the direct encoding schemes like QAOA, which require $m$ qubits—one per variable—this approach achieves polynomial compression in the qubit count.
-The desired solution of the combinatorial problem is now encoded in terms of a non-classical Hamiltonian.
-The compression ratio depends on the polynomial order $k$, allowing us to tackle larger MaxCut instances with fewer qubits.
-All methods are evaluated under a unified encoding framework with $k=2$ to ensure fairness.
-To evaluate solution quality across varying scales, we conducted experiments with 1,200 measurement shots, averaging results over multiple trials.
-As shown in Figure 5(Left), scaling sVQNHE parameters—through enhancements in the NN component or other optimizations—consistently widens its performance gap over competing methods.
-Specifically, sVQNHE, which employs NN-guided hierarchical optimization and NN post-processing, significantly outperforms sign-VQE and brickwork-VQE.
-Sign-VQE adopts the final PQC structure from sVQNHE (a two-layer sign ansatz with two diagonal and two non-diagonal layers) but omits the NN-guided optimization and post-processing used in sVQNHE.
-Brickwork-VQE, in contrast, utilizes a VQE with a brickwork ansatz, as detailed in [36].
-Figure 5(Left) reveals that, without NN guidance and post-processing (sign-VQE vs.
-brickwork-VQE), the sign ansatz shows no clear advantage over the brickwork ansatz, with comparable solution quality and similar measurement costs.
-However, sVQNHE’s hybrid approach - integrating hierarchical optimization guided by NN and NN post-processing to Sign Ansatz - yields a marked improvement.
-For instance, the performance gap between sVQNHE and baseline methods increases by approximately $6\%\approx(0.96-0.90)/0.90$ for 9-vertex graphs and $19\%\approx(1-0.84)/0.84$ for 45-vertex graphs when using parameter set 2.
-Additionally, Figure 5(Right) shows that sVQNHE significantly reduces single-step measurement consumption compared to both sign-VQE and brickwork-VQE.
-With parameter set 2, it achieves $86\%\approx(72-10)/72$ and $85\%\approx(306-46)/306$ improvements in quantum resource efficiency for 9- and 45-vertex systems, respectively.
-Additionally, in the Supplementary Information, we provide a preliminary comparison of resource consumption for larger instances, further highlighting sVQNHE’s advantages in scalability and efficiency under large-scale conditions.
-This underscores its potential for large-scale quantum applications.
+For sVQNHE, coefficients of variation after stabilization are $cv_{1}=0.01618,cv_{2}=0.015912$ (a) and $cv_{1}=0.01534,cv_{2}=0.01487$ (b), indicating reduced variability in the second layer.
+Under realistic device noise (Figure 4(b)), sVQNHE again achieves faster convergence to near-ground-state energies and exhibits lower variability in the refined second layer ($CV_{1}=0.01534\to CV_{2}=0.01487$).
+This demonstrates that the hybrid quantum-neural architecture possesses an intrinsic error-mitigating capability.
+The classical NN module can be continuously tuned during noisy optimization to partially compensate for circuit-induced errors, as previously analyzed in related variational hybrid frameworks [zhang2023variational].
+These noisy experiments highlight that sVQNHE not only maintains its advantages in ideal settings but also exhibits superior robustness and reliability under the combined effects of shot noise and incoherent errors.
+4. Toward Near-Term Quantum Utility in Large-Scale Combinatorial Optimization
+Finally, we investigate whether the sVQNHE framework can deliver meaningful quantum utility in practically relevant combinatorial optimization problems using highly compressed qubit encodings.
+We consider two canonical NP-hard problems:
+    MaxCut Problem (Erdős–Rényi graphs, up to 1485 vertices),
+    Maximum Clique Problem (Erdős–Rényi graphs, 135 vertices).
+Both problems are encoded using Pauli Correlation Encodings (PCE) [sciorilli2025towards], achieving polynomial qubit compression: $O(n^{k})$ classical variables are mapped to only $n$ qubits with $k=2$ (MaxClique) or $k=4$ (large MaxCut).
+For MaxCut, under noisy sampling conditions, as shown in Figure 5 (a), sVQNHE consistently outperforms the baselines in solution quality.
+The performance gap widens with scale: for $m=9$, sVQNHE achieves $\sim 6\%$ improvement over baselines (e.g., 0.96 vs.
+0.90), rising to $\sim 19\%$ for $m=45$ (1.00 vs.
+0.84) using optimized parameter set 2 (enhanced NN depth and learning rate).
+Without NN components, sign-VQE and brickwork-VQE perform comparably, indicating that the sign ansatz alone offers no inherent edge over brickwork; the gains stem from sVQNHE’s hybrid integration.
+Figure 5 (b) highlights resource efficiency: sVQNHE reduces single-step measurement costs by $\sim 86\%$ for $m=9$ (10 vs.
+72) and $\sim 85\%$ for $m=45$ (46 vs.
+306) compared to baselines.
+This efficiency arises from NN-guided layer-wise optimization, which minimizes the number of gradient evaluations.
+Preliminary results in Supplementary Information extend this to larger instances ($m>45$), confirming sVQNHE’s scalability under noisy sampling.
+For large-scale instances, we use state-vector simulations on graphs with $m=1485$ vertices and edge probability $0.01$, encoded via quartic PCE ($k=4$), requiring $n=12$ qubits.
+This setup tests the case where direct encodings are infeasible in the NISQ and early FTQC eras.
+We generate five random instances and compare sVQNHE against: (1) Free-Energy Machine [shen2025free] (FEM), a physics-inspired classical solver minimizing variational free energies via gradient descent; (2) D-Wave’s Simulated Annealing Sampler [dwave_ocean_sdk] (SA); and (3) a pure NN solver, using only the classical components of the corresponding sVQNHE method, without the quantum part.
+For each instance, as in the sampling case, we compute the cut value for all methods and normalize relative to the maximum cut across the four solvers.
+Averaging over the five graphs yields average relative solution qualities: sVQNHE (0.9989), FEM (0.9989), NN (0.9800), and SA (0.8066) (see Figure 5 (c)).
+sVQNHE and FEM deliver near-optimal solutions, with sVQNHE slightly outperforming FEM by exploiting quantum correlations for finer exploration of the rugged landscape.
+The pure NN lags due to underparameterization in high-dimensional spaces, while SA struggles with sparse graphs (low edge probability), as annealing heuristics favor denser structures [johnson1989optimization, myklebust2015solving, hamerly2019experimental].
+For MaxClique, we evaluate on graphs with $m=135$ vertices and edge probability $p=0.5$.
+Encoding uses PCE with $k=2$, compressing to $10$ qubits.
+Methods include sVQNHE, its pure real NN component (identical classical part without quantum module), VQNHE (hardware-efficient ansatz with equivalent number of quantum parameters), greedy (degree-based iterative addition), and FEM.
+Non-greedy methods require post-processing (thresholding + greedy repair) to ensure valid cliques, as continuous relaxations may yield infeasible assignments.
+Over 3 instances, average relative solution qualities, normalized to the maximum across methods) rank as: sign-VQNHE ($1.00$), Greedy ($0.89$), NN and VQNHE ($0.81$), FEM ($0.26$) (see Figure 5 (d)).
+sVQNHE outperforms greedy’s typical $maxclique\approx\log_{2}135$ (though greedy requires no post-processing, its local search limits global exploration).
+The superior performance of sVQNHE highlights the critical value of its quantum module.
+Compared to the pure NN baseline (which uses the exact same classical neural network component but lacks the quantum phase-learning layers), sVQNHE delivers significantly higher solution quality, demonstrating that the quantum part (diagonal gates $W_{i}$ for phase structure learning) actively contributes meaningful computational power rather than merely consuming resources.
+This advantage is further validated against VQNHE (with equivalent parameter count): sVQNHE’s innovative co-learning procedure of phase and amplitude components with adaptive layer-wise growth harnesses quantum resources more effectively, avoiding barren plateaus and reducing measurement costs.
+In addition, although methods relying on continuous energy landscape optimization face greater challenges in this rugged landscape, relative to the continuous optimization method FEM, which claims generality for COPs without Ising reductions [shen2025free], sVQNHE offers superior global search in constraint-intensive, rugged landscapes.
+In contrast, FEM’s mean-field independence assumption ($\prod_{i}P_{i}(\sigma_{i})$) fails to capture the strong all-pair connectivity constraints inherent to cliques, leading to diffuse marginal distributions and higher post-processing losses [jagota1995approximating, shen2025free].
+This underscores sVQNHE’s enhanced universality for combinatorial optimization problems.
+These scaling experiments on MaxCut and Maximum Clique demonstrate that sVQNHE consistently delivers superior or competitive performance across diverse combinatorial optimization landscapes.
+On MaxCut, a problem with relatively smoother energy surfaces, sVQNHE achieves near-optimal solutions comparable to state-of-the-art classical solvers like FEM while offering slight advantages in rugged exploration via quantum correlations, together with substantial reductions in measurement overhead through NN-guided optimization.
+On Maximum Clique, a paradigmatically rugged, constraint-intensive problem with strong all-pair correlations, sVQNHE markedly outperforms all baselines and validates the critical value of its quantum module.
+Overall, sVQNHE emerges as a robust, scalable hybrid framework that leverages qubit-efficient encodings and synergistic quantum-classical computation to address both smooth and rugged combinatorial optimization problems effectively, positioning it as a promising candidate for near-term quantum utility in combinatorial optimization.
 Figure 5: Scaling test of sVQNHE using the MaxCut problem as an example.
-(Left) The relative solution quality for the MaxCut problem, comparing parameter sets 1 and 2 across different problem sizes.
+(a) The relative solution quality for the MaxCut problem, comparing parameter sets 1 and 2 across different problem sizes.
 The methods sign-VQNHE, brickwork-VQE, and sign-VQE correspond to the sVQNHE method with a 2-layer sign ansatz, the VQE method with brickwork ansatz, and the VQE method with 2-layer sign ansatz, respectively.
-The depth of brickwork ansatz is determined according to the reference [36], that is, the circuit depth is sublinearly related to the number of vertices in the graph.
+The depth of brickwork ansatz is determined according to the reference [sciorilli2025towards], that is, the circuit depth is sublinearly related to the number of vertices in the graph.
 For sign-VQE, the ansatz is the same as the circuit finally generated by the corresponding sVQNHE scheme finally generated, utilizing a two-layer sign ansatz with two diagonal and two non-diagonal layers.
 The value range of relative solution quality is $[0,1]$.
-The closer to 1, the better the performance is compared with other algorithms (see Equation (14) in the Method section for details).
+The closer to 1, the better the performance is compared with other algorithms (see Equation (LABEL:eq:R_e) in the Method section for details).
 For example, in the case of 45 vertices, the quality of the solution of sVQNHE is relatively optimal regardless of whether the parameter set is 1 or 2.
-(Right) The single-step measurement consumption for the same problem sizes.
+(b) The single-step measurement consumption for the same problem sizes.
 Parameter set 2 features an increased number of parameters in the NN component and doubles the number of iterations compared to parameter set 1.
 The circuit depth in the brickwork architecture scales sub-linearly with the number of vertices.
+(c) Exact state-vector simulation for MaxCut on Erdős–Rényi graphs with $m=1485$ vertices.
+Gray error bars represent one standard deviation.
+(d) Exact state-vector simulation for MaxClique on Erdős–Rényi graphs with $m=135$ vertices.
+Gray error bars represent one standard deviation.
 Discussion
-We present sVQNHE, a hybrid quantum-classical framework that strategically optimizes the roles of quantum and classical resources.
-Compared to the original VQNHE, our approach significantly improves measurement efficiency, substantially suppresses the numerical instability of the neural network, and enhances overall convergence.
-Notably, sVQNHE leverages commuting diagonal gates to encode quantum sign structures, allowing efficient gradient estimation with far fewer measurements.
-The classical neural network is tasked with learning wavefunction amplitudes and guiding parameter updates, effectively sidestepping the barren plateau issue commonly seen in deep quantum circuits.
-This division of labor enables a layer-wise optimization strategy that remains stable even under noise and sampling limitations.
-Our numerical simulations demonstrate compelling improvements across a broad class of computational problems.
-In the 6-qubit J1-J2 model, the sVQNHE method with a single-layer assumption achieves more than 7-fold the average accuracy and 19-fold the convergence speed of the 2-layer hardware-efficient VQE.
-Compared with the baseline NN, the average accuracy is improved by nearly two orders of magnitude.
-In sign-free models such as the TFIM and Ising chains, sVQNHE consistently outperforms QAOA and VQE, achieving broader expressivity and better scalability.
-For combinatorial optimization, such as MaxCut on 45-node graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$ compared to brickwork ansatz-based VQE.
-Taken together, these results highlight sVQNHE’s broad applicability, efficiency, and scalability.
-By tightly integrating neural networks with quantum circuits along with a revised role that each component plays in the new algorithm, sVQNHE offers a practical path toward quantum utility on NISQ and early fault-tolerant devices.
-Looking ahead, this framework can be further extended by incorporating advanced machine learning models and better measurement strategies, and applied to a wider class of many-body simulation and large-scale optimization problems.
-Extending sVQNHE to tackle PQC-based quantum machine learning is also conceptually straightforward.
-As quantum hardware continues to evolve, we expect sVQNHE to remain a strong candidate for bridging theoretical capabilities with real-world impact.
+The results presented in this work support a broader computational principle: hybrid quantum–classical wavefunction learning becomes significantly more scalable when amplitude and phase are trained by different resources rather than by a single quantum circuit.
+Recent efforts in variational quantum computing have focused on improving expressivity, ansatz structure, or optimizer robustness while still relying on the quantum circuit to learn both components of the wavefunction.
+Our findings suggest this paradigm may be unnecessarily burdensome for near-term hardware.
+Demonstrated across a broad range of applications in this study, when the amplitude landscape is learned classically, and the quantum circuit is dedicated solely to learning the phase (sign) structure, optimization becomes measurably easier, gradients remain accessible at larger system sizes, and noise sensitivity decreases without sacrificing expressivity, as evidenced by sVQNHE’s benign variational landscape and gradual transfer mechanism.
+Importantly, the performance improvements observed across quantum simulation and combinatorial optimization arise from the division of labor itself.
+Amplitude learning thrives under classical updates, whereas phase learning benefits from quantum interference in shallow circuits.
+This explains why sVQNHE’s scaling laws differ from VQE or QAOA: gradient variance no longer collapses exponentially, and measurement cost does not inflate with circuit depth or Hamiltonian complexity.
+In practice, this enables deeper phase-layer stacking on noisy hardware, with inherent error mitigation under realistic gate noise and finite sampling.
+Empirically, in frustrated many-body problems like the 6-qubit J1-J2 model, sVQNHE reduces mean absolute energy error by $98.9\%$ and variance by $99.6\%$ relative to neural baselines, converging $\sim 19\times$ faster than hardware-efficient VQE.
+Similar gains hold for the 9-qubit 2D Heisenberg model and even sign-free systems like the 9-qubit TFIM or 12-qubit Ising, with error reductions up to $81.1\%$.
+Extending to molecular systems, sVQNHE yields superior ground-state energies and fidelities for a H2O molecule across geometries.
+For combinatorial optimization, it excels in MaxCut on Erdős–Rényi graphs, improving solution quality by up to $19\%$ and reducing measurements by $\sim 85\%$ in sampling-based tests up to 45 vertices.
+More broadly, these findings shift how hybrid algorithms are conceptualized: the neural network as the primary learning module, while quantum hardware supplying essential but targeted phase transformations.
+This implies scaling quantum advantage via architectures where quantum roles are small yet critical, as seen in sVQNHE matching the classical Free-Energy Machine on a 1,485-vertex MaxCut instance and a 135-vertex MaxClique instance.
+Overall, explicit quantum phase learning paired with classical amplitude modeling constitutes a principled hybrid strategy.
+By refining the quantum hardware’s role, this work reinforces that quantum utility may depend more on complementary classical-quantum designs than on emulating universal circuits.
 Methods
@@
 Theorem 1
-(Expressiveness of the Quantum Circuit Ansatz)[40, 41, 12, 42, 43] Let $\mathfrak{g_{1}}$ and $\mathfrak{g_{2}}$ be the Lie algebras corresponding to the generators $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, where:
+(Expressiveness of the Quantum Circuit Ansatz)[larocca2023theory, allcock2024dynamical, holmes2022connecting, kazi2024analyzing, kiani2020learning] Let $\mathfrak{g_{1}}$ and $\mathfrak{g_{2}}$ be the Lie algebras corresponding to the generators $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, where:
 $\mathcal{G}_{1}=\{Z_{D_{2j}}\mid j=1,2,\ldots,\lfloor\frac{m}{2}\rfloor\}\bigcup\{Y_{i}\mid i=1,2,\ldots,n\},$
@@
 Then, any matrix commuting with $M$ must share its block structure.
+Theorem 2
+(Benign Landscape of sVQNHE) For an $l$-layer ($l\geq 1$) sVQNHE algorithm with a commuting diagonal ansatz $W_{l}(\vec{\theta})$, the cost function $\mathcal{L}(\vec{\theta})=\langle\psi(\vec{\theta})|H|\psi(\vec{\theta})\rangle$, where $\psi(\vec{\theta})=Z_{f}F_{l}W_{l}(\vec{\theta})|\phi\rangle=W_{l}(\vec{\theta})|\phi_{f}\rangle$, exhibits at worst polynomially vanishing gradient variance $\text{Var}[\partial_{\theta_{k}}\mathcal{L}]\in\Omega(1/\text{poly}(n))$ if the support set size of $|\phi_{f}\rangle$ is polynomially bounded in $n$.
+Here, $Z_{f}$ is the normalization coefficient and $|\phi_{f}\rangle=Z_{f}F|\phi\rangle$.
+Proof: To prove the theorem, we establish that the variance of the gradient $\partial_{\theta_{k}}\mathcal{L}$ with respect to a parameter $\theta_{k}$ in the $l$-th layer ansatz $W_{l}(\vec{\theta})$ scales at worst polynomially with the number of qubits $n$, avoiding exponential vanishing[cerezo2021cost], due to polynomial constrained input and the commuting diagonal structure.
+Since $W_{l}(\vec{\theta})$ is a diagonal unitary composed of commuting Z-related gates, it can be expressed as
+$W_{l}(\vec{\theta})=\sum_{x\in\{0,1\}^{n}}e^{i\xi(x,\vec{\theta})}|x\rangle\langle x|,$
+where $\xi(x,\vec{\theta})=\sum_{k=1}^{K}\theta_{k}g_{k}(x)$ for basis functions $g_{k}(x)$ (e.g.
+for the parameters $\theta_{k}$ of gate $R_{zz}$ applied on the $k_{1},k_{2}$ qubit, $g_{k}(y)=(-1)^{y_{k_{1}}+y_{k_{2}}}$, where $y=|y_{1}y_{2}\ldots y_{n}\rangle$).
+The gradient is:
+$\partial_{\theta_{k}}\mathcal{L}=2\operatorname{Re}\left[\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger}(\vec{\theta}))HW_{l}(\vec{\theta})|\phi_{f}\rangle\right],$
+where
+$\partial_{\theta_{k}}W_{l}(\vec{\theta})=i\sum_{x}g_{k}(x)e^{i\xi(x,\vec{\theta})}|x\rangle\langle x|,$
+so
+$\partial_{\theta_{k}}W_{l}^{\dagger}(\vec{\theta})=-i\sum_{x}g_{k}(x)e^{-i\xi(x,\vec{\theta})}|x\rangle\langle x|.$
+To assess barren plateaus, compute the variance over random initializations $\vec{\theta}\sim[0,2\pi)^{K}$:
+$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]=\mathbb{E}_{\vec{\theta}}[(\partial_{\theta_{k}}\mathcal{L})^{2}]-(\mathbb{E}_{\vec{\theta}}[\partial_{\theta_{k}}\mathcal{L}])^{2}.$
+The mean vanishes: $\mathbb{E}_{\vec{\theta}}[\partial_{\theta_{k}}\mathcal{L}]=0$, as $\mathbb{E}_{\vec{\theta}}[e^{i\xi(x,\vec{\theta})}]=0$ for non-constant $\xi$ due to phase cancellation over $0,2\pi)$.
+Thus, the variance is:
+$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]=\mathbb{E}_{\vec{\theta}}[(\partial_{\theta_{k}}\mathcal{L})^{2}].$
+Substitute the gradient
+$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]=\mathbb{E}_{\vec{\theta}}[\left[\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger})HW_{l}|\phi_{f}\rangle\right]^{2}+\left[\langle\phi_{f}|W_{l}^{\dagger}H(\partial_{\theta_{k}}W_{l})|\phi_{f}\rangle\right]^{2}+2\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger})HW_{l}|\phi_{f}\rangle\langle\phi_{f}|W_{l}^{\dagger}H(\partial_{\theta_{k}}W_{l})|\phi_{f}\rangle].$
+Now, substitute the expressions:
+$W_{l}(\vec{\theta})=\sum_{x}e^{i\xi(x,\vec{\theta})}|x\rangle\langle x|,\quad\partial_{\theta_{k}}W_{l}(\vec{\theta})=i\sum_{y}g_{k}(y)e^{i\xi(y,\vec{\theta})}|y\rangle\langle y|,\quad\partial_{\theta_{k}}W_{l}^{\dagger}(\vec{\theta})=-i\sum_{y}g_{k}(y)e^{-i\xi(y,\vec{\theta})}|y\rangle\langle y|.$
+Then:
+$\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger})HW_{l}|\phi_{f}\rangle=-i\sum_{y,z}g_{k}(y)e^{-i\xi(y,\vec{\theta})}\langle\phi_{f}|y\rangle\langle y|H|z\rangle e^{i\xi(z,\vec{\theta})}\langle z|\phi_{f}\rangle,$
+$\langle\phi_{f}|W_{l}^{\dagger}H(\partial_{\theta_{k}}W_{l})|\phi_{f}\rangle=i\sum_{y,z}g_{k}(z)e^{-i\xi(y,\vec{\theta})}\langle\phi_{f}|y\rangle\langle y|H|z\rangle e^{i\xi(z,\vec{\theta})}\langle z|\phi_{f}\rangle.$
+and
+$\langle\phi_{f}|(\partial_{\theta_{k}}V^{\dagger})HV|\phi_{f}\rangle\cdot\langle\phi_{f}|V^{\dagger}H(\partial_{\theta_{k}}V)|\phi_{f}\rangle$
+$=\sum_{y_{1},z_{1},y_{2},z_{2}}g_{k}(y_{1})g_{k}(z_{2})e^{-i(\xi(y_{1})-\xi(z_{1}))}e^{i(\xi(y_{2})-\xi(z_{2}))}\langle\phi_{f}|y_{1}\rangle\langle y_{1}|H|z_{1}\rangle\langle\phi_{f}|y_{2}\rangle\langle y_{2}|H|z_{2}\rangle\langle z_{1}|\phi_{f}\rangle\langle z_{2}|\phi_{f}\rangle.$
+Averaging over $\vec{\theta}$:
+$\mathbb{E}_{\vec{\theta}}\left[e^{-i(\xi(y_{1})-\xi(z_{1})+\xi(y_{2})-\xi(z_{2}))}\right]=\delta_{y_{1}z_{1}}\delta_{y_{2}z_{2}},$
+because the expectation over $\vec{\theta}$ cancels unless the phase differences are zero (e.g., $y_{1}=z_{1}=y_{2}=z_{2}$).
+The same argument applies to other items in $\text{Var}[\partial_{\theta_{k}}\mathcal{L}]$.
+Thus we have
+$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]\geq\Omega(\sum_{y}g_{k}^{2}(y)|\langle y|H|y\rangle|^{2}|\langle\phi_{f}|y\rangle|^{4}).$
+For the form with restricted support set size, where $|\langle\phi_{f}|y\rangle|>0$ only at $poly(n)$ positions—instead of the type where $2^{n}$, by the Cauchy–Schwarz inequality we have:
+$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]\geq\Omega(\frac{1}{poly(n)}).$
+[…]
 Gradual transfer mechanism
@@
 The parameter shift rule is defined as follows:
-$\gradient_{\theta}f(x;\theta)=\frac{1}{2}[f(x;\theta+\frac{\pi}{2})-f(x;\theta-\frac{\pi}{2})].$ (8)
+$\nabla_{\theta}f(x;\theta)=\frac{1}{2}[f(x;\theta+\frac{\pi}{2})-f(x;\theta-\frac{\pi}{2})].$ (8)
 To maximize the efficiency of executing the parameter shift rule, we exploit the commutative property of $W(\theta^{W})$.
@@
 $\frac{\partial E_{P_{i}}(\theta^{W})}{\partial\theta_{k}^{W}}=\frac{1}{2}(E_{P_{i}}(\theta_{k}^{W}+\frac{\pi}{2})-E_{P_{i}}(\theta_{k}^{W}-\frac{\pi}{2})).$ (9)
-Since $W_{k}$ and $F$ are all commutative, and $E_{P_{i}}(\theta^{W})=\bra{\psi_{0}}W^{\dagger}(\theta^{W})F^{\dagger}P_{i}FW(\theta^{W})\ket{\psi_{0}}$, we have
+[…]
 $E_{P_{i}}(\theta_{k}^{W}\pm\frac{\pi}{2})$
-$=\bra{\psi_{0}}W^{\dagger}(\theta^{W})F^{\dagger}W_{k}^{\dagger}(\pm\frac{\pi}{2})P_{i}W_{k}(\pm\frac{\pi}{2})FW(\theta^{W})\ket{\psi_{0}}.$ (10)
+[…]
 Equation (9) and (Measurement Strategy and Cost) imply that to obtain the gradient of $\theta^{W}$, we only need to add the measurement of Puali string $L_{i}=W_{k}^{\dagger}(\frac{\pi}{2})P_{i}W_{k}(\frac{\pi}{2})$.
@@
 In summary, due to the commutation properties of the Sign Ansatz, our algorithm requires fewer measurements compared to general algorithms that rely on parameterized quantum circuits.
-For instance, considering the 1D J1-J2 model presented below (11), if $|p|\geq 2$ in the Sign Ansatz, the total number of measurements in a single iteration of our algorithm is equal to $13(n_{1}+n_{2})+$ 1, where $n_{1}$ is the number of $\langle i,j\rangle$, and $n_{2}$ is the number of $\langle\langle i,j\rangle\rangle$.
+For instance, considering the 1D J1-J2 model presented below (11), if $|p|\geq 2$ in the Sign Ansatz, the total number of measurements in a single iteration of our algorithm is equal to $3(n_{1}+n_{2})+1$, where $n_{1}$ is the number of $\langle i,j\rangle$, and $n_{2}$ is the number of $\langle\langle i,j\rangle\rangle$.
 $H=J_{1}\sum_{\langle i,j\rangle}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{j}^{z})+J_{2}\sum_{\langle\langle i,j\rangle\rangle}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y}+\Delta_{2}\sigma_{i}^{z}\sigma_{j}^{z})+B_{H}\sum_{i}\sigma_{i}^{z}.$ (11)
@@
 The details of our numerical simulation
-We use TensorCircuit[48] to efficiently obtain all the quantum circuits and the neural network numerical results with the following Ansatz setting, and the details of the VQNHE algorithm are in the TensorCircuit Demo.
-[…]
-The Sign Ansatz design equivalents for the sampling and noise versions of the numerical results with finite shots are detailed as follows.
-A combination of $YM$, $YP$, $C_{z}$, and $R_{z}$ gates equivalently replaces the $R_{zz}$ gate.
-Similarly, the $R_{y}$ gate is equivalently replaced by a combination of $XP$, $XM$, and $R_{z}$ gates.
-Here, $YM$ and $YP$ denote positive and negative rotations by $\frac{\pi}{4}$ around the Y-axis of the Bloch sphere, respectively, while $XM$ and $XP$ denote positive and negative rotations by $\frac{\pi}{4}$ around the X-axis of the Bloch sphere, respectively.
-The two-qubit gates used in our cases connect the edges of the model (1D J1-J2 model and 2D Heisenberg model).
-[…]
-In the design of Hardware Efficient Ansatz, where the $CNOT$ is only connecting the two nearest qubits (2D Heisenberg Model), n is the number of repetitions.
-Note that after repeating the above block n times, one more layer of $R_{y}$ and one more layer of $R_{z}$ are added.
-For example, for HEA2 used in this paper, there are three $R_{y}$ layers, three $R_{z}$ layers, and two $CNOT$ layers.
-Similarly, the $CNOT$ gate in the numerical simulation of the finite-shot version is equivalently replaced by a combination of $YM$, $YP$, and $C_{z}$ gates.
-Metrics
-In the experiment, we use the following metrics to evaluate the performance of the method:
-  • Mean absolute error ratio $R_{MAE}$:
-    $R_{MAE}=\frac{MAE_{1}-MAE_{0}}{MAE_{0}}.$ (12)
-    The metric $R_{MAE}$ represents the relative reduction in mean absolute error (MAE) of different algorithms compared to the baseline algorithm (NN) with the same number of iterations.
-Here $MAE_{0}$ represents the MAE of the baseline algorithm (NN) being compared.
-The smaller the $R_{MAE}$ value, the better the compared algorithm is relative to NN.
-  • Variance ratio $R_{Var}$:
-    $R_{Var}=\frac{Var_{1}-Var_{0}}{Var_{0}}.$ (13)
-    The metric $R_{Var}$ represents the relative reduction in the variance of different algorithms compared to the baseline algorithm (NN) given the same number of iterations.
-Among them, $Var_{0}$ represents the variance of the baseline algorithm (NN) being compared.
-The smaller the $R_{Var}$ value, the more stable the compared algorithm is relative to NN, and the smaller the variability of the results.
-  • Estimated approximation ratio $R_{e}$ for the MaxCut problem:
-    $R_{e}=\frac{MaxCut_{algo}}{MaxCut_{max}}.$ (14)
-    The metric $R_{e}$ provides an estimated approximation ratio based on the best-known result for the MaxCut problem, which measures the relative solution quality.
-Here, $MaxCut_{max}$ represents the maximum MaxCut value achieved across all evaluated algorithms, while $MaxCut_{algo}$ denotes the MaxCut value obtained by the specific algorithm under assessment.
-A value of $R_{e}$ approaching $100\%$ indicates superior performance relative to the other algorithms compared.
-Acknowledgements
-CYH acknowledges support from National Ke Research and Development Program of China (2024FA1306400), Natural Science Foundation of China (22373085).
-AH gratefully acknowledges the sponsorship from Research Grants Council of the Hong Kong Special Administrative Region, China (Project No.
-CityU 11200120), City University of Hong Kong (Project No.
-7005615, 7006103), CityU Seed Fund in Microelectronics (Project No.
-9229135), and Hon Hai Research Institute (Project No.
-9231594).
-YYS acknowledges support from the National Natural Science Foundation of China (Grant No.
-12404575), the Shandong Provincial Natural Science Foundation (Grant No.
-ZR2022LLZ008), the Natural Science Foundation of Shanghai (Grant No.
-23ZR1469600), and the Shanghai Sailing Program (Grant No.
-23YF1452600).
-Code and data availability
-The source code and data associated with this project are publicly available. They can be accessed at the following GitHub repository: https://github.com/renmengzhen/sVQNHE
-Supplementary
-Optimizing Quantum Resources with Sign Ansatz VQNHE
-In this section, we evaluate the quantum resource efficiency of the sVQNHE framework against the Hardware Efficient Ansatz VQE.
-First, we test different ansatz on the 9-qubit 2D Heisenberg model with the initial state $\ket{++...++}$.
-The results are shown in Figure 6, it can be seen that sVQNHE uses very few quantum resources to converge to the desired accuracy compared to both neural networks and VQE, which implies that the sVQNHE reaches the desired accuracy with fewer quantum resources compared to VQE in the number of circuits, and the number of parameters.
-Besides, since the zero quantum resource results indicate that without the assistance of Sign Ansatz the neural network itself cannot reach good accuracy, we can secondly claim that this framework can greatly improve the expressiveness of the neural network with limited quantum resources.
-In addition, we also verified the quantum resource efficiency of sVQNHE relative to VQE on the J1-J2 model.
-The results are shown in the Table 3.
-Figure 6: 9-qubit Heisenberg Model Result.
-The x-axis shows sorted results by the converged energy with different random initial parameters.
-The y-axis is the energy difference from the ground state.
-The red, green, and blue dashed lines are responding to Hardware Efficient Ansatz with repetitions equal to 1, 2, and 3, and the solid line is Complex MLP(3,3) with Sign Ansatz, where the total number of parameters in the neural network is 317.
-Both of them start from the initial state $\ket{+...+}$.
-The dotted line $99.99\%$ Energy is 0.0001 times the ground state energy.
-Table 3: Convergence Efficiency Comparison for the 6 qubit J1-J2 Model
-Method Average Optimization Steps Success Probability (%) Summary
-sVQNHE 132.67 60 faster convergence with higher reliability
-VQE (2-layer HEA) 1676.00 5 more steps and has a lower success rate
-Explanation: This table compares the convergence efficiency of sVQNHE and VQE (with a 2-layer hardware-efficient ansatz) for the 6-qubit J1-J2 model.
-The metrics include the average number of optimization steps required to reach 99.45% of the ground state energy and the probability of success.
-sVQNHE outperforms VQE by requiring fewer steps (132.67 vs.
-1676) and achieving a higher success rate (60% vs.
-5%), demonstrating superior efficiency in quantum resource utilization.
-Sampling Efficiency Verification
-This section presents a compelling argument for performing sampling from the quantum circuit component rather than the classical neural network component that describes the amplitude.
-Our objective is to verify the computational advantages in the main text argument.
-As illustrated in Figure 7, we compare the convergence curve of our model with that shown in Figures 3 and 4 of the main text.
-The key difference is that the set of bitstrings $s$ here is obtained by sampling from the classical neural network component.
-Compared to the results in Figures 3 and 4 of the main text, we observe that performing sampling from the quantum circuit component is better; fewer shots can yield more stable iterations and faster convergence.
-This suggests that our quantum circuit can effectively sample from complex probability distributions that are intractable for classical computers, thereby providing a high-quality solution space that closer approximates the true solution space.
-Figure 7: Sampling Efficiency Verification under 2D Heisenberg Hamiltonian.
-The figure illustrates the results obtained under varying numbers of qubits and shots.
-Sampling is conducted from the classical neural network component.
-The x-axis denotes the number of iterations, while the y-axis represents the energy.
-The solid dotted line indicates the ground state energy value of the corresponding model.
-Preliminary Resource Consumption Analysis for Larger Instances
-Table 4: Resource Comparison for Larger Instances
-Qubits Encoding Vertices
-brickwork-VQE
-measurement cost
-Comparison
-sVQNHE
-measurement cost
-17 2 408 3000 $>$ 409
-3 2040 24150 $>$ 4081
-30 2 1305 9612 $>$ 1306
-3 12180 141246 $>$ 24361
-Explanation: This table compares the measurement costs of sVQNHE and brickwork-VQE under 2nd and 3rd order polynomial encodings for 17 and 30 qubits.
-The vertex count represents the maximum number of graph vertices for the maxcut problem corresponding to each qubit count and encoding scheme.
-Measurement cost reflects the number of measurement circuits required per iteration.
-sVQNHE consistently requires fewer circuits, demonstrating superior efficiency in quantum resource utilization.
-In this section, we compare the measurement requirements for a single iteration of the brickwork-VQE and sVQNHE schemes under 2nd and 3rd order polynomial encodings[36], with equal shots per operator.
-As shown in the Table.
-4, the savings in measurements, defined as the reduction when using sVQNHE compared to brickwork-VQE, scale linearly with the number of vertices in the graph.
-While this scaling is linear, the absolute savings become significantly large for graphs with many vertices, resulting in a substantial efficiency advantage for sVQNHE in large-scale quantum systems.
+We use TensorCircuit[zhang2023tensorcircuit] to efficiently obtain all the quantum circuits and the neural network numerical results with the following Ansatz setting, and the details of the VQNHE algorithm are in the TensorCircuit Demo.